### v1

# Generative Regression Based Watch Time Prediction for Video Recommendation: Model and Performance

Abstract.
Watch time prediction (WTP) has emerged as a pivotal task in short video recommendation systems, designed to encapsulate user interests. Predicting users’ watch times on videos often encounters challenges, including wide value ranges and imbalanced data distributions, which can lead to significant bias when directly regressing watch time. Recent studies have tried to tackle these issues by converting the continuous watch time estimation into an ordinal classification task. While these methods are somewhat effective, they exhibit notable limitations. For instance, the discretization process often hinges heavily on bucket partitioning, which inherently compromises prediction accuracy. Moreover, the relationships existing between various partition intervals are seldom fully exploited, thereby ignoring the potential for effective error correction.
Inspired by language modeling, we propose a novel Generative Regression (GR) paradigm for WTP based on sequence generation. This approach employs structural discretization to enable the lossless reconstruction of original values while maintaining prediction fidelity. By formulating the prediction problem as a numerical-to-sequence mapping, and with meticulously designed vocabulary and label encodings, each watch time is transformed into a sequence of tokens. To expedite model training, we introduce the curriculum learning with an embedding mixup strategy which can mitigate training-and-inference inconsistency associated with teacher forcing. We evaluate our method against state-of-the-art approaches on four public datasets and one industrial dataset. We also perform online A/B testing on Kuaishou, a leading video app with about 400 million DAUs, to demonstrate the real-world efficacy of our method. The results conclusively show that GR outperforms existing techniques significantly. Furthermore, we successfully apply GR to another regression task in recommendation systems, i.e., Lifetime Value (LTV) prediction, which highlights its potential as a novel and effective solution to general regression challenges.
Recommendation, Watch-time prediction, Generative regression
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The Web Conference 2025; April 28–May 2, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems
1. Introduction
Figure 1. Predictive paradigm comparison among ordinal regression methods CREAD (a) and TPM (b), and our generative regression (c). Red lines indicate the discretization structure.
In recent years, online short video content has a remarkable surge with the rapid development of short video social media platforms such as TikTok and Kuaishou, which spurs efforts to optimize recommendation systems for streaming players (Covington et al., 2016; Davidson et al., 2010; Liu et al., 2019, 2021). Unlike traditional Video on Demand (VOD) platforms such as Netflix and Hulu, short video platforms in scrolling mode automatically play content without the user clicking action for desirable video choice, rendering traditional metrics such as click-through rates obsolete (Gao et al., 2022b; Gong et al., 2022). Under these circumstances, the watch time of videos has emerged as a critical metric for measuring user engagement and experience (Covington et al., 2016; Wang et al., 2020; Wu et al., 2018; Yi et al., 2014). Continuous video watching means users’ immersion and enjoyment of the platform, enhancing the probability of further user retention and conversion. Consequently, accurate watch time estimation enables platforms to recommend videos prolonging users’ viewing, which impacts key business metrics such as Daily Active Users (DAU) and drives revenue growth.
In contrast to limited and discrete actions such as liking, following, and sharing, watch time generally exhibits a wide range and long-tailed distribution. This characteristic renders it fundamentally a regression problem for prediction. However, the continuous nature of watch time also makes it vulnerable to outliers and potential predictive bias, thereby increasing modeling complexity and challenges of accurate prediction. Some methods (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a) optimize watch time prediction from a debiasing perspective but have not yet adequately addressed the core challenges of regression. Some others (Sun et al., 2024; Lin et al., 2023) transform the prediction problem into an Ordinal Regression (OR) task by employing a series of binary classifications across various predefined time intervals, as separately shown in Fig. 1(a) and Fig. 1(b). Nevertheless, such a modeling paradigm still exhibits two major limitations as follows:
Firstly, conditional dependencies among time intervals are not fully leveraged. The dependencies are solely reflected in the definition of the labels, predictions across different time intervals are often produced independently, thereby hindering the potential for effective error correction.
Secondly, the ordinal regression paradigm decomposes the regression target into a series of binary classification problems defined by fixed time intervals, which inherently confines the labeling of samples within these rigid divisions. As a result, model performance is highly contingent on the method of time interval segmentation. Coarse segmentation, i.e., wider spans, may overlook subtle differences, while overly fine segmentation, i.e., narrower spans, can result in insufficient data, degrading the model’s ability to learn effective features (Sun et al., 2024).
In response to these limitations above, inspired by the recent success of Large Language Models (LLMs) (Touvron et al., 2023; Brown, 2020; Zhao et al., 2023b), we propose a novel universal regression paradigm, called Generative Regression (GR), which effectively utilizes dependencies among multi-step predictions and does not strictly rely on fixed time interval divisions. GR addresses the issues above as follows:
On the one hand, as shown in Fig. 1(c), the complete watch time prediction task is decomposed into a sequential generation task, where each step predicts a part of the total watch time. The output of each time step serves as input for the next one, thereby constituting a conditional and sequential modeling process. The objective is to predict a sequence of time slots, whose sum constitutes the continuous regression target. This generative regression paradigm not only ingeniously inherits the advantage of previous ordinal regression methods (Sun et al., 2024; Lin et al., 2023; Frank and Hall, 2001; Li and Lin, 2006) by decomposing the regression task into multi-classification subtasks to simplify the prediction process, but also leverages dependencies between steps to accurately and progressively approximate the total watch time.
On the other hand, unlike ordinal regression methods (Sun et al., 2024; Lin et al., 2023) that restrict outputs to binary classification within fixed time intervals, our GR model provides the flexibility of allowing each predictive step to choose from a vocabulary of tokens, where each token represents a distinct time slot in positive real number space. This enlarged solution space enables GR to generate a broader set of potential sequences, thereby improving its capacity to generalize across diverse watching behaviors and leading to more accurate and personalized predictions. A comprehensive comparison of GR with other regression methods is shown in Tab. 1.
For token definition and watch time segmentation, we propose a data-driven unified vocabulary construction method, which mitigates token imbalance and eliminates reliance on manual design, and a label encoding strategy allows a lossless restoration of watch time values, thereby enhancing the model’s generality and generalization capability. To accelerate model convergence, we adopt curriculum learning (Bengio et al., 2009) strategy that combines teacher forcing (Venkatraman et al., 2015) with standard training to alleviate training-and-inference inconsistency, commonly known as exposure bias (Venkatraman et al., 2015; Ding and Soricut, 2017). Besides, leveraging our insights into the training process, we propose an embedding mixup method to compensate for output-to-input gradients. This approach enhances model performance at a lower computational cost by leveraging the semantic additivity of tokens while ensuring consistency between training and inference.
Table 1. A comprehensive comparison with existing deep regression methods. ‘H’, ‘M’, and ‘L’ indicate high, medium, and low respectively. Complexity represents computation overhead during inference, with $m$, $n$, $d$, and $t$ indicating the last hidden feature size, discretization segments, vocabulary size, and sequence length, respectively.
Methods Predictive Paradigm Discre -tization Accuracy Versatility Complexity
VR Discriminative - L H $\mathcal{O}(m)$
TPM (Lin et al., 2023) Discriminative Tree-based M M $\mathcal{O}(mn)$
CREAD (Sun et al., 2024) Discriminative Linear M M $\mathcal{O}(mn)$
GR Generative Sequential H H $\mathcal{O}(mdt)$
The contributions of this paper are as follows:
  1. (i)
    We introduce a novel generation framework for predicting watch time, which inherits the benefits of structured discretization and adeptly utilizes interval relationships for the progressive and precise estimation of total watch time.
  2. (ii)
    To enhance the generality and adaptability of GR, we develop a data-driven unified vocabulary design and a label encoding method. Additionally, we introduce curriculum learning with embedding mixup to mitigate exposure bias and accelerate model training.
  3. (iii)
    Extensive online and offline experiments demonstrate that GR significantly outperforms existing SOTA models. We further analyze the impact of key factors like vocabulary design and model architecture, to provide a clear understanding of the mechanisms and processes underlying GR.
  4. (iv)
    Last but not least, we successfully apply GR to another regression task in recommendation systems, Lifetime Value (LTV) prediction, which indicates its potential as a novel and effective solution to general regression challenges.
2. Related Work
Watch time prediction aims to estimate the video watch time based on the user’s profile, historical interactions, and video characteristics. Value regression (VR) directly predicts the absolute value of watch time, assessing model accuracy by mean square error (MSE). WLR (Covington et al., 2016) applies watch duration as weights within its loss function, yet it may suffer from significant bias due to its weighting mechanism. D2Q (Zhan et al., 2022) segments data by duration, predicting watch time quantiles within these segments to alleviate duration bias. CWM (Zhao et al., 2024) introduces a counterfactual watch time, estimating a video’s hypothetical full watch time to gauge user interest. D2Co (Zhao et al., 2023a) differentiates actual user interest from duration bias and noisy watching using a duration-wise Gaussian mixture model. However, these methods have not yet adequately addressed the core challenges of regression. Some solutions transform the problem into ordinal regression. CREAD (Sun et al., 2024) introduces an error-adaptive discretization technique to construct dynamic time intervals. TPM (Lin et al., 2023) utilizes hierarchical labels to model relationships across varying granularities of time intervals. Yet, these approaches are unable to fully capitalize on the interdependencies among these intervals and heavily rely on time interval segmentation.
Ordinal regression, also known as ordinal classification, is a type of predictive modeling strategy employed when the outcome variable is ordinal and the relative order of labels is important, such as age prediction (Niu et al., 2016), monocular depth perception (Fu et al., 2018), and head-pose estimation (Hsu et al., 2018). CNNOR (Liu et al., 2017) introduces a deep learning model for ordinal regression tailored to small datasets. SORD (Diaz and Marathe, 2019) employs a soft ordinal label approach to train the ordinal regression model. Additionally, (Li et al., 2021) crafts a probabilistic embedding strategy specifically for ordinal regression tasks. Ordinal regression has not been applied to watch time prediction until the introduction of CREAD (Sun et al., 2024) and TPM (Lin et al., 2023). These models follow the traditional approach of decomposing regression into multiple binary classification tasks, achieving significant benefits.
Figure 2. The framework of GR model, which adopts an encoder-decoder RNN-based architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner and employs the curriculum learning with embedding mixup (CLME) strategy to alleviate training-and-inference inconsistency introduced by teacher forcing.
Sequence generation requires the model to understand the context of an input sequence and generate an appropriate output sequence. Early researches focus on natural language processing (NLP) tasks such as machine translation (Sutskever, 2014; Elman, 1990; Graves and Graves, 2012; Cho, 2014), text summarization (Bahdanau, 2014; Vaswani, 2017), and dialogue generation (Devlin, 2018; Raffel et al., 2020). In recommendation systems, sequential recommendation methods have been proposed to capture sequential patterns. GRU4Rec (Hidasi, 2015) is a session-based recommendation model with GRU. SASRec (Kang and McAuley, 2018) utilizes a self-attention mechanism to capture both long-term and short-term user preferences. BERT4Rec (Sun et al., 2019) employs a bidirectional transformer to encode item sequences. However, sequence generation models in recommendation systems predominantly focus on predicting the sequence of user behaviors, and their application to watch time prediction remains unexplored.
3. Method
3.1. Problem Formulation
Given a dataset $\mathcal{D}=\left\{(\bm{u_{i}},\bm{v_{i}},y_{i})\right\}_{i=1}^{N}$, where $\bm{u_{i}}$ and $\bm{v_{i}}$ represent the user-side features (such as user ID, static profile, and historical behaviors etc.) and the item-side (videos in this paper) features (e.g. tags, duration and category) of the $i$-th example respectively  [^1] We omit the context-side features for simplicity., $y_{i}\in\mathbb{R}$ is the corresponding watch time of the $i$-th example collected from recommendation system logs. Value regression (VR) methods aim to learn a function $f(\cdot)$ that directly maps the input features to a real-valued output, i.e., $y_{i}=f([\bm{u_{i}};\bm{v_{i}}])$.
Sequence generation in GR is mostly based on autoregressive language modeling. Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$, where each element $w_{j}$ represents a predefined time slot (e.g. 5 seconds, 10 seconds, etc.). The details of vocabulary construction are presented in Sec. 3.3. Here, these time slots are analogous to tokens in language models (LMs). Thus, “token” and “time slot” will be used interchangeably in the sequel. The vocabulary embedding matrix is denoted as $\bm{E}\in\mathbb{R}^{V\times D}$, where $V$ is the vocabulary size, and $D$ is the dimension of the time slot embeddings.
We decompose $y_{i}$ into a sequence of tokens $\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}$, where $s_{i}^{t}\in\mathcal{V}$ and $T_{i}$ is the length of the sequence. We call this process label encoding, which is detailed in Sec. 3.4. On the other hand, we design a label decoding function $\text{g}(\cdot)$ such that derives the original watch time $y_{i}$ from $\bm{s_{i}}$, i.e., $y_{i}=\text{g}(\bm{s_{i}})=\sum_{t=1}^{T_{i}}\text{g}(s_{i}^{t})\in\mathbb{R}$. Our goal is to train a sequence generation model, given the user and video features $(\bm{u_{i}},\bm{v_{i}})$, generates the corresponding sequence of watch time slots $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, and in turn, from which the predicted watch time $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$ approximates the actual watch time $y_{i}$.
3.2. The Generative Regression (GR) Model
As shown in Fig. 2, GR adopts an encoder-decoder RNN-based architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner.
3.2.1. Encoder
Unlike traditional sequence-to-sequence tasks or user behavior modeling in recommendation systems, watch time prediction does not inherently depend on history action sequences. Consequently, directly using an RNN as the encoder — designed for processing temporally dependent sequential data — is inappropriate and may result in suboptimal feature representation.
To ensure the generality and simplicity of the model, we use a feed-forward network (FFN) as the encoder. However, it can be replaced with any complex model architecture. Formally, the encoder extracts user and video features to produce a fixed-length hidden feature $\bm{h_{i}}\in\mathbb{R}^{1\times D}$ that will be fed to the decoder as follows:
(1) $\bm{h_{i}}=\bm{W_{L}}\cdot(...\text{relu}(\bm{W_{2}}\cdot(\text{relu}(\bm{W_{1}}\cdot\bm{x}_{i}))))$
where $\bm{x_{i}}=[\bm{u_{i}};\bm{v_{i}}]$, $\bm{W_{1}},...,\bm{W_{L}}$ are weight parameters of FFN.
3.2.2. Decoder
[…]
As illustrated in Fig.2, the decoder generates the sequence of watch time slots $\bm{\hat{s_{i}}}=\{\hat{s}_{i}^{1},...,\hat{s}_{i}^{j},...,\hat{s}_{i}^{T_{i}}\}$ conditioned on the encoder output $\bm{h_{i}}$ and the preceding subsequence. Specifically, at time step $t$ in training, the output token $\hat{s}_{i}^{t}$ will be computed as
(2) $\hat{s}_{i}^{t}=\arg\max_{w\in\mathcal{V}}P_{\theta}(w\mid\bm{h_{i}},\bm{\hat{s}_{i}^{<t}})$
where $\theta$ is the model parameter and $\bm{\hat{s}_{i}^{<t}}$ is the tokens generated before. Utilizing the chain rule, the overall probability of generating the sequence can be expressed as
(3) $P_{\theta}(\bm{s_{i}}\mid\bm{h_{i}})=P_{\theta}(s_{i}^{1},...,s_{i}^{T_{i}}\mid\bm{h_{i}})=\prod_{t=1}^{T}P(s_{i}^{t}\mid\bm{h_{i}},\bm{{s}_{i}^{<t}})$
There remain three issues to be addressed: how to construct an effective vocabulary? how to encode $y_{i}$ into a sequence $\bm{s_{i}}$? and how to optimize the model? which are detailed in the following sections.
3.3. Vocabulary Construction
As mentioned before, tokens in vocabulary $\mathcal{V}$ represent predefined watch time slots that enable the model to generate sequences closely approximating the actual watch time values. Based on our cognition of the deep regression task, three principles are designed to guide the construction of vocabulary.
  • Completeness: The vocabulary $\mathcal{V}$ must be able to represent all watch time values $\{y_{i}\}_{i=1}^{N}$ using a finite number of tokens almost without loss [^2] The label space is continuous, it is impossible to recover the original watch time from a discretized sequence with zero error.. Also, each token must be unique.
  • Balance: The frequencies of tokens should be relatively uniform to prevent class imbalance.
  • Adaptability: The vocabulary should remain consistent to ensure scalability and adaptability across various datasets.
One intuitive strategy is to select watch time values from the dataset as tokens based on several fixed percentiles, yet failing to meet the completeness principle. An alternative is to select watch time values as tokens based on one fixed percentile, then subtract the token values from all watch time values that exceed them, repeating this process until the residuals become negligible, which fails to meet the balance principle.
Algorithm 1 Constructing Vocabulary with dynamic percentiles
1:Dataset labels $\bm{Y}=\{y_{j}\}_{j=1}^{N}$, initially empty Vocabulary $\mathcal{V}=\{\}$, precision threshold $\epsilon$, start percentile $q_{\text{start}}$, end percentile $q_{\text{end}}$, decay rate $\alpha$, minimal restoration error $\epsilon$.
2:Sort $\bm{Y}$ in descending order to obtain $\hat{\bm{Y}}$.
3:Initialize iteration counter $i=1$, error metric $err=\infty$, current percentile $q=q_{\text{start}}$
4:while $err>\epsilon$ do
5:     Compute the $q$-percentile $z_{i}$ of $\bm{Y}$.
6:     if $z_{i}=0$ then $\triangleright$ Terminate if the percentile value is zero
7:         break
8:     end if
9:     Insert $z_{i}$ into vocabulary $\mathcal{V}$.
10:     Update $\hat{\bm{Y}}$ using:
$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if }\hat{y}_{j}<z_{i},\\ \hat{y}_{j}-z_{i},&\text{otherwise}\end{cases}$
11:     Update the error metric $err$: $err=\max\{\frac{\hat{y}_{j}}{y_{j}}\}_{j=1}^{N}$
12:     Update percentile $q$ with decay rate $\alpha$: $q=\max(q\cdot\alpha,~{}q_{\text{end}})$
13:     Increase $i$: $i=i+1$.
14:end while
15:return $\mathcal{V}$
To meet these principles above, we develop a data-drived vocabulary construction algorithm based on dynamic percentile, detailed in Algorithm. 1. Specifically, the algorithm starts with an initially large percentile $q_{start}$ and decreases it according to the decay rate $\alpha$, ensuring that the percentile does not fall below the endpoint $q_{end}$. This strategy expedites the reduction of tail values, rapidly decreasing the variance among updated values, which effectively mitigates the challenges posed by the long-tailed distribution in the dataset, for which we provide detailed validation in the experimental section.
Under the generative regression framework, it is important to note that the process of constructing a vocabulary and converting the task target (e.g. watch time) into a sequential token format (a.k.a label encoding) resembles the syntax construction of a language. However, we do not contend that the aforementioned strategy is the most optimal one.
3.4. Label Encoding
Given the vocabulary $\mathcal{V}$, we need to do label encoding, i.e., transforming the watch time values $\{y_{i}\}_{i=1}^{N}$ to corresponding target sequences $\{\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}\}_{i=1}^{N}$. To facilitate the learning process with the encoded labels (i.e., transformed token sequences), we also offer three principles for label encoding as follows:
  • Correctness: $y_{i}=\sum_{t=0}^{T_{i}}w_{i}^{t}+\epsilon$, where $\epsilon$ is a small value (e.g. 0.1% relative error) which can be omitted.
  • Minimal sequence length: $T_{i}$ is the possible minimal length of the token sequence selected to satisfy correctness for $y_{i}$.
  • Monotonicity: $s_{i}^{1}\geq s_{i}^{2}\geq...\geq s_{i}^{T_{i}}$ is required.
The minimum sequence length requirement simplifies the learning process. Additionally, the non-decreasing order captures the dynamic changes in user interest while watching videos, making the sequence generation process progressively (from easy to hard). Without loss of generality, the token indices in the vocabulary are also sorted in descending numerical order.
To this end, we employ a greedy strategy for the shortest sequence length, starting from the largest possible watch time slot and decreasing progressively, decomposing the total watch time $y_{i}$ into a sequence of watch time slots.
3.5. Optimization and Inference
3.5.1. Vanilla Training Process
Similar to language modeling, the model predicts the next token $s^{t}$ using the ground truth tokens $s^{<t}$. The learning objective is to maximize the negative log-likelihood of the generated sequence over the training set and effectively minimize the difference between the predicted sequence and the ground truth sequence, which is computed by the cross-entropy loss as follows:
(4) $\mathcal{L}_{ce}=-\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log P_{\theta}(\hat{s}_{i}^{t}\mid\bm{h_{i}},\hat{s}_{i}^{<t})$
Following previous works (Lin et al., 2023; Sun et al., 2024), we employ the huber loss (Huber, 1992) to guide regression:
(5) $\mathcal{L}_{huber}=\mathcal{L}_{\delta}(y_{i},\hat{y}_{i})=\begin{cases}\frac{1}{2}(y_{i}-\hat{y}_{i})^{2}&\text{if}|y_{i}-\hat{y}_{i}|\leq\delta,\\ \delta\cdot(|y_{i}-\hat{y}_{i}|-\frac{1}{2}\delta)&\text{otherwise}\end{cases}$
where $\hat{y_{i}}=\sum_{t=1}^{T_{i}}g(\hat{s}_{i}^{t})$, $\delta$ acts as a threshold, toggling between quadratic and linear losses to balance sensitivity and robustness against outliers. Therefore, the loss function we need to optimize is
(6) $\mathcal{L}=\mathcal{L}_{ce}+\lambda\cdot\mathcal{L}_{huber}$
$\lambda$ is a hyperparameter that balances the two losses. To improve model efficiency, we adopt a teacher forcing (TF) strategy, which directly feeds the ground truth output $s_{i}^{t}$ as input at step $t+1$ to guide model training. However, since the ground truth is unknown during inference, the discrepancy in input for the decoder leads to the well-known exposure bias problem, which can degrade model performance (Goodman et al., 2020).
3.5.2. Curriculum Learning with Embedding Mixup (CLEM)
We introduce a Curriculum Learning (CL) strategy that combines TF and vanilla training (a.k.a free running mode) by progressively incorporating stochasticity into the training process. Specifically, to predict $\hat{s}_{i}^{t}$, we randomly choose ground truth token $s_{i}^{t-1}$ with a dynamic probability $p$ as the sampling rate.
To warm up, we start with $p\approx 1$, indicating that the model predominantly relies on the ground truth tokens. We then adjust the probability $p$ using a non-linear decay strategy, which increases the likelihood of sampling from the predicted sequence. This enables the model to gradually adapt to the inference stage. Formally,
(7) $p=p_{0}\cdot\frac{\omega}{\omega+e^{\left(\frac{\tau}{\omega}\right)}}$
where $\tau$ is the training iteration and $\omega>0$ influences the shape of the descent curve to ensure a seamless transition from higher to lower values. This strategy addresses exposure bias by learning to predict with both ground truth and previous prediction as input. In Sec. 4.5, we also conduct detailed experimental comparisons of additional strategies such as linear and exponential decay.
(a) Watch time embeddings visualization during training.
(b) Probability difference score among tokens during training.
Figure 3. Watch time embedding with a weighted sum of token embeddings (left) and the probability distribution difference among tokens for each $\hat{s}_{i}^{t}$ (right). Best viewed in color.
We notice that GR can capture the inherent relationship between different tokens. As the cardinality of $\mathcal{V}$ is not as large as that in LMs, it is hard to visualize $\mathcal{V}$’s embedding matrix $\bm{E}$. As an alternative, we visualize the target numerical embeddings of $y_{i}$ by $\bm{e_{i}}=\sum_{t=1}^{T_{i}}r_{t}\bm{E}[\hat{s_{i}}^{t},:]$, $r_{t}=\frac{g(\hat{s}_{i}^{t})}{y_{i}}$ is the contribution of $\hat{s}_{i}^{t}$ to the target value $y_{i}$. As shown in Fig. 3(a), values beginning with the same token $w_{j}$ are clustered, and the initial tokens of similar numeric values $\text{g}(w_{j})$ are also relatively close to each other. This clustering feature indicates a coherent structure in the embedding space, suggesting that the model effectively captures similarities among the tokens.
Although this is what we expect the model to learn, it can also lead to the issue of confusion among similar tokens. In other words, the model may struggle to effectively distinguish between the target token and its neighboring tokens. As noted in Sec. 3.4, tokens are arranged in non-increasing order within the vocabulary $\text{g}(w_{1})>\text{g}(w_{2})>...>\text{g}(w_{V})$. We illustrate this phenomenon in Fig. 3(b) by computing the averaged probability difference of each token relative to its neighbors. The results indicate that the predicted probability of the target token is closely aligned with the surrounding tokens, which contributes to an increase in the final prediction error. Since our token sequences are encoded in non-increasing order, it becomes challenging for subsequent tokens to bridge the gap between $\hat{y}_{i}$ and ground truth $y_{i}$ once the first token, denoted as $\hat{s_{i}}^{1}$ is wrong. Additionally, the argmax operation in Eq. (2) will disrupt the gradient propagation path, which further complicates the learning process.
We observe that employing a curriculum training strategy with inference by using a beam search ensemble significantly boosts model performance. However, it incurs more computational cost. Given the previously discussed properties of tokens and sequences, the numerical differences between adjacent tokens are minimal. To address this, we propose to integrate the embedding sequences of the preceding tokens through an ensemble approach during the training process to improve the prediction of the next token.
Embedding Mixup (EM). The cohort is centered on the current predicted token $\hat{s}^{t}_{i}$ with window size $n_{w}$, the mixup region is $[\hat{s}^{t}_{i}-b,\hat{s}^{t}_{i}+b]$ and $b=\lfloor\frac{n_{w}}{2}\rfloor$. We have
(8) $\bm{z_{i}}^{t}=\sum\limits_{j=0}^{n_{w}}\sigma_{j}\cdot\bm{E}[\hat{s}^{t}_{i}+j-b,:]$
(9) $\sigma_{j}=\frac{exp(-\rho_{j})}{\sum_{k=0}^{n_{w}}exp(\rho_{\hat{s}^{t}_{i}+k-b})}$
where $\sigma_{j}$ recalculates the fusion weights of tokens in the fixed window size, $\rho_{j}$ is the logit predicted by the decoder at step $t$. And $\bm{z_{i}}^{t}$ will replace original $\bm{E}[\hat{s}^{t}_{i},:]$ as the input for the next step. This re-weighted mixup approach leverages the numerical semantics and additivity property inherent in our tokens. The re-weighting ensures that the semantic space is aligned, eliminating any discrepancies in scale. EM offers three benefits: 1) It reduces the learning complexity of the model by merging representations of tokens within a fixed window, thereby preventing significant errors; 2) The integration leverages the predicted scores from the previous steps, enhancing the information transfer from output to input in recurrence structure; and 3) it ensures consistency between training and inference while lowering inference cost.
3.5.3. Inference Process
During inference, the encoder extracts $\bm{h_{i}}$ from input features $[\bm{u_{i}},\bm{v_{i}}]$, the decoder begins with the ¡SOS¿ token, and sequentially generates the prediction sequence $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$. The process continues until the token ¡EOS¿ is generated, which signifies the completion of the sequence. Finally, the predicted watch time is computed as $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$.
4. Experiments
This section presents extensive experiments to demonstrate the effectiveness of the GR model. Five research questions are explored in these experiments:
  • RQ1: How does GR compare to state-of-the-art methods in terms of prediction accuracy of watch time?
  • RQ2: What is the effect of vocabulary design on the performance of GR and why?
  • RQ3: How do various neural network architectures perform within the GR framework?
  • RQ4: What impact does CLEM have on the GR model, and how do different training strategies affect performance?
  • RQ5: How does GR perform on other regression tasks?
Due to the space limit, additional ablation study results are available at A.2.
4.1. Experiment Settings
4.1.1. Datasets
Two publicly available datasets and one industrial dataset are used to evaluate the proposed method. The large-scale industrial dataset (Indust for short) is sourced from a real-world streaming short-video app Kuaishou with about 400 million DAUs and multi-billion impressions each day. We collect interaction logs for 4 days and utilize the subsequent day’s data for evaluation. The CIKM16333https://competitions.codalab.org/competitions/11161, sourced from the CIKM16 Cup competition, is designed to predict user engagement duration in online search sessions. It contains 310,302 sessions, 122,991 items, and an average session length of 3.981. KuaiRec (Gao et al., 2022a) is a real-world dataset collected from the video view logs of the Kuaishou app, comprising data from 7,176 users, 10,728 items, and 12,530,806 impressions. In addition to directly predicting watch time, some researchers (Lin et al., 2023) focus on predicting the watch ratio, which can be used in conjunction with video duration to calculate watch time. Therefore, we report both watch time and watch ratio results on the KuaiRec dataset.
Table 2. Performance comparison among different approaches on KuaiRec, CIKM16 and Indust.
Method KuaiRec (watch time) KuaiRec (watch ratio) CIKM16 Indust
MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$
VR 7.634 0.504 - 0.385 0.691 - 1.039 0.641 - 46.343 0.588
WLR (Covington et al., 2016) 6.047 0.525 4.167% 0.375 0.698 1.013% 0.998 0.672 4.836% - -
D2Q (Zhan et al., 2022) 5.426 0.565 12.103% 0.371 0.712 3.039% 0.899 0.661 3.120% - -
TPM (Lin et al., 2023) 3.456 0.571 13.294% 0.361 0.734 6.223% 0.850 0.676 5.460% 41.486 0.593
CREAD (Sun et al., 2024) 3.307 0.594 17.857% 0.369 0.738 6.802% 0.865 0.678 5.772% 39.979 0.597
GR (ours) 3.196 0.614 21.825% 0.333 0.753 8.972% 0.815 0.691 7.80% 38.528 0.604
    Here, the best and second best results are marked in bold and underline, respectively. $\uparrow$ indicates that the higher the value is, the better the performance is, while $\downarrow$ signifies the opposite. Each experiment is repeated 5 times and the average is reported.
Table 3. Performance gain on online A/B testing.
A/B test APP usage time Average app usage per user
+0.055% +0.072%
APP launch counts Video consumption time
+0.042% +0.129%
4.1.2. Compared Methods
Considering baseline methods compared in prior studies (Sun et al., 2024; Lin et al., 2023), we compare several state-of-the-art methods (Covington et al., 2016; Zhan et al., 2022; Sun et al., 2024; Lin et al., 2023) with our GR. For a fair comparison, similar model structures except for the output layers and corresponding loss functions are employed for these approaches. Details of the compared methods are provided in ”watch time prediction” of Sec. 2.
4.1.3. Metrics
To evaluate the proposed method, we follow previous work (Lin et al., 2023; Sun et al., 2024) and utilize two performance metrics:
  • Mean Average Error (MAE): It quantifies regression accuracy by averaging the absolute deviations between predicted values $\{\hat{y_{i}}\}_{i=1}^{N}$ and actual values $\{y_{i}\}_{i=1}^{N}$ as follows:
    (10) $MAE=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$
  • XAUC (Zhan et al., 2022): This measure assesses the concordance between the predicted and actual ordering of watch time values. We uniformly sample pairs from the test set and calculate the XAUC by determining the percentile of samples that are correctly ordered. A higher XAUC indicates better model performance.
4.2. Performance Comparison (RQ1)
4.2.1. Offline
We compare the proposed method GR with 5 existing models in terms of watch time and watch ratio, results are presented in Tab. 2. Specifically, for watch time prediction, GR exhibits a 4.117% relative decrease in MAE and a 1.917% improvement in XAUC on the CIKM16 dataset. On the KuaiRec benchmark, significant improvements are achieved, with a 3.356% reduction in MAE and a 3.367% increase in XAUC relative to the second-best method. As for Indust dataset, GR exhibits a 3.629% relative decrease in MAE and a 1.001% improvement in XAUC compared to the CREAD. Regarding watch ratio predictions, all models exhibit significant performance gain due to the elimination of video duration bias. Despite this, GR maintains the best performance, boasting a 7.756% reduction in MAE and a 2.033% improvement in XAUC. The comprehensive improvements in both MAE and XAUC substantiate the superiority of the GR method. To confirm that the observed enhancements in the GR model are not solely attributable to increased model parameters, further comparisons with parameter-equivalent models can be found in the appendix A.2.1.
4.2.2. Online A/B Testing
Besides offline experiments, we also conduct an online A/B test on Kuaishou App to demonstrate the real-world efficacy of our method. Considering that Kuaishou serves about 400 million users daily, doing experiments from 2.5% of traffic involves a huge population of more than 10 million users, which can yield highly reliable results. Most recommendation systems follow a two-stage framework: initially retrieving a set of candidate items and then selecting the top-ranked items in the ranking phase. Watch time prediction, a component of the ranking stage, plays a crucial role. Items with higher predicted watch times are more likely to be recommended.
The online experiment have been launched on the system for six days and the results are in Tab. 3. The metrics for online experiments include app usage time, average app usage per user, app launch counts and video consumption time (accumulated watch time). The results demonstrate that GR consistently boosts performance in watch time related metrics, with an improvement by 0.055% on app usage time, 0.072% on average app usage per user, and significant 0.129% on video consumption time, substantiating its potential to significantly enhance real-world user experiences.
4.3. Vocabulary Construction Analysis (RQ2)
Here we examine the effect of the vocabulary construction method. Besides the proposed Dynamic Quantile algorithm, two commonly used methods are considered: Manual that designs the vocabulary based on experience, e.g., using values like 1ms, 3ms and 5ms, then scaling them by 10, 100, and so on until exceeding the maximum watch time in the dataset. Binary starts with the smallest unit of watching duration, i.e., milliseconds, as the first token, with each subsequent token being twice the value of its predecessor until exceeding the maximum watch time in the dataset. Tab. 4 presents the experimental results. We can see that the proposed dynamic quantile method outperforms the other two strategies. Notably, our method is nearly automatic, which makes it more efficient than the manual and binary vocabulary construction methods.
We further analyze token frequency distribution, i.e., counting the occurrences of each token in the vocabulary, and results are shown in Fig. 4. We sort all tokens in descending order according to frequencies and select the top 15 for analysis and comparison. We can see that in the binary method, nearly half of the tokens are scarcely used, while the manual method exhibits a highly imbalanced distribution. In contrast, our dynamic quantile method achieves a more balanced distribution, further validating the efficacy of the proposed algorithm.
Table 4. Comparison of vocabulary construction methods.
Vocabulary design KuaiRec CIKM16
[…]
Manual 3.281 0.604 0.825 0.685
Binary 3.268 0.605 0.821 0.687
Dynamic quantile 3.196 0.614 0.815 0.691
4.4. Generative Architecture Analysis (RQ3)
To investigate the impact of different network architectures on prediction performance, we compare three popular sequence modeling structures: RNN, GRU, and LSTM, by substituting the decoder structure in the GR model. Results are presented in Fig. 5, from which we can see that GR is not dependent on a specific network architecture and achieves desirable outcomes across various architectures. Among them, LSTM exhibits the best performance, likely because GR ultimately sums all predicted tokens to estimate watch time, resulting in strong dependencies between tokens in the sequence. LSTM is particularly adept at capturing these dependencies across different token positions.
Figure 4. Token distribution comparison among vocabulary construction methods: (a) Manual, (b) Binary, (c) Dynamic.
4.5. Effect of Curriculum Learning with Embedding Mixup (RQ4)
Here, we investigate the impact of the proposed curriculum learning (CL) with embedding mixup (EM) strategy by designing several variants to perform comparisons from three perspectives.
  • To study the contribution of CLEM, we design three ablation variants. w/o CLEM uses free running mode to directly feed $\bm{E}[\hat{s}_{i}^{t-1},:]$ to generate $\hat{s}_{i}^{t-1}$. EM with TF indicates embedding mixup strategy with full teacher forcing, i.e., with fixed sampling rate $p$ = 1. CL w/o EM removes embedding mixup strategy.
  • To explore the effect of different dynamic decay strategy, we design linear decay and exponential decay instead of Eq. (7), thereby constructing two variants linear and exponential.
  • To examine the effect of the curriculum strategy and free running, we construct two model variants with fixed sampling rates, where $p$ is set to 0.5 and 0, respectively.
Table 5. Ablation study on the strategy of curriculum learning with embedding mixup.
Method KuaiRec CIKM16
MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
(a) GR 3.196 0.614 0.815 0.691
(b) w/o CLEM 3.416 0.584 0.858 0.674
(c) EM with TF 3.241 0.604 0.844 0.684
(d) CL w/o EM 3.359 0.588 0.849 0.679
(e) linear 3.205 0.613 0.818 0.690
(f) exponential 3.211 0.613 0.819 0.690
(g) ${p=0.5}$ 3.208 0.612 0.820 0.690
(h) ${p=0}$ 3.283 0.593 0.846 0.681
The results in Tab. 5 indicate that both strategies in CLEM enhance performance. Specifically, comparing row (a) with row (c), we can see that CL increases XAUC by 1.656% and reduces MAE by 1.38% on the KuaiRec dataset, while comparing row (a) with row (d) shows EM achieves even greater improvements, with 4.235% increase in XAUC and 4.853% reduction in MAE. Removing all the strategies results in substantial performance degradation for the GR model, with a loss of 5.137% in XAUC and 6.884% in MAE.
The sampling rate decay coefficients significantly impact both metrics. The proposed curriculum strategy achieves a gain of 2.65% in MAE and 3.42% in XAUC on KuaiRec, comparing row (a) with row (h). Although different nonlinear decay strategies yield similar results in terms of XAUC, they still improve MAE. These findings indicate that the CLEM strategy improves the model’s accuracy of watch time prediction.
Figure 5. Performance comparison of different generative architectures on (a) KuaiRec and (b) CIKM16.
4.6. Performance on LTV Prediction Task (RQ5)
GR is a versatile regression framework. To explore its potential in other tasks, we conduct additional experiments on the Lifetime Value (LTV) prediction task. Following the settings  (Weng et al., 2024), we evaluate GR on Criteo-SSC [^4] https://ailab.criteo.com/criteo-sponsored-search-conversion-log-dataset/ and Kaggle [^5] https://www.kaggle.com/c/acquire-valued-shoppers-challenge, using MAE and Spearman rank correlation (Spearman’s $\rho$) as performance metrics. We consider all baseline methods compared in prior studies (Weng et al., 2024).
Table 6. Performance comparison on LTV datasets.
Method Criteo-SSC Kaggle
[…]
Two-stage (Drachen et al., 2018) 21.719 0.2386 74.782 0.4313
MTL-MSE (Ma et al., 2018) 21.190 0.2478 74.065 0.4329
ZILN (Wang et al., 2019) 20.880 0.2434 72.528 0.5239
MDME (Li et al., 2022) 16.598 0.2269 72.900 0.5163
MDAN (Liu et al., 2024) 20.030 0.2470 73.940 0.4367
OptDist (Weng et al., 2024) 15.784 0.2505 70.929 0.5249
GR (ours) 12.996 0.3026 67.035 0.5334
Tab. 6 presents the results of various methods in LTV prediction. These results show that our approach significantly outperforms the previous works across ordinal and numeric performance metrics, achieving 17.66% improvement of MAE and 20.79% improvement on Spearman’s $\rho$ on Criteo-SSC, compared with the state-of-the-art method OptDist (Weng et al., 2024). Given that the compared methods are specifically optimized for the LTV task, these findings demonstrate the robustness and extensibility of GR, underscoring its potential for application to other continuous value prediction tasks within the domain of recommendation systems.
5. Conclusion
This paper proposes a novel regression paradigm Generative Regression (GR) to accurately predict watch time, which addresses two key issues associated with existing methods. First, many ordinal regression methods struggle to accurately recover durations due to discretization, with performance heavily reliant on the chosen time binning strategy. Second, while ordinal regression implicitly constrains the probability distribution along the estimation path to exhibit a decreasing trend, existing methods have not fully leveraged this property. The generative regression approach builds upon autoregressive modeling and offers a promising exploration space. We also introduce embedding mixups and curriculum learning during training to accelerate model convergence and enhance performance. Extensive offline and online experiments show that GR significantly outperforms the state-of-the-art models. Additionally, we investigate its applicability to lifetime value (LTV) prediction, where our generative regression method also surpasses the SOTA models, highlighting its potential as a novel and effective solution to general regression challenges.
Appendix A Appendices
A.1. Experimental Settings
Unless otherwise stated, we adopt the following training protocol. The hyperparameters $\lambda$ in equation Eq. (6) are deliberately set to 10. Our GR architecture comprises a two-layer Multi-Layer Perceptron (MLP) for the encoder and a single-layer Long Short-Term Memory (LSTM) network for the decoder. We employ a dropout rate of 0.3 to prevent overfitting and set the learning rate to 5e-4. As for vocabulary construction in Algorithm. 1, the decay rate $\alpha$ is set to 0.99, precision threshold $\epsilon$ is set to 1e-6, start percentile $q_{start}$ and end percentile $q_{e}nd$ are respectively set to 0.9999 and 0.90. The $\omega$ in Eq. (7) is set to 17000. In Eq. (8), the window size $n_{w}$ is set to half the length of the vocabulary. We train for 20 epochs in total with a batch size of 1024, using the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ to minimize the loss function.
Table 7. The impact of model parameters on the performance between different approaches on KuaiRec and CIKM16.
Method Parameters KuaiRec CIKM16
MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
VR 0.86M 7.634 0.504 1.039 0.641
TPM 0.86M 3.456 0.571 0.850 0.676
CREAD 0.86M 3.307 0.594 0.865 0.678
VR-large 4.34M 7.556 0.521 0.998 0.651
TPM-large 4.34M 3.432 0.577 0.843 0.682
CREAD-large 4.34M 3.293 0.599 0.857 0.680
GR (ours) 4.18M 3.196 0.614 0.815 0.691
A.2. Additional Ablation Study Results
A.2.1. Model Parameters
To ascertain that the enhancement observed in the GR model is not merely a consequence of an increase in model parameters, we conducted a comparative analysis, as presented in Tab. 7. Specifically, we standardized the parameters of the SOTA models, CREAD (Sun et al., 2024) and TPM (Lin et al., 2023), to a uniform level. The results indicate that, when evaluated with an equivalent parameter scale, the GR model consistently surpasses the previous SOTA methods across a range of metrics.
A.2.2. Hyperparameter for Vocabulary Construction
We conduct an ablation study to investigate the effects of the decay rate $\alpha$ and end percentile $q_{end}$ in Algorithm. 1 on the performance of GR. All experiments maintain a constant start percentile $q_{start}$ set at 0.9999 and the results are present in Tab. 8. Under the same end percentile $q_{end}$, models with a decay rate of 0.99 consistently outperform those with a decay rate of 0.98. Within identical settings for $\alpha$, the configurations with $q_{end}=0.9$ yield the best performance results, followed by $q_{end}=0.85$.
Figure 6. (a) Dynamic probability across different epochs for varying parameter $\omega$. The curves show the decrease in sampling rate over time. (b) The impact of $\omega$ on MAE and XAUC for the KuaiRec dataset.
A.2.3. Hyperparameter for Curriculum Learning (CL)
Here, we check the effects of the hyperparameters $\omega$ in Eq. (7) which influences the shape of descent curve to ensure a seamless transition from higher to lower values. As shown in Fig. 6(a), four distinct values of $\omega$, i.e., 4600, 8500, 13000, and 17000, are chosen to control the dynamic probability, gradually reducing it to zero at the 5th, 10th, 15th, and 20th epoch, respectively. The results are presented in Fig. 6(b). We can see that $\omega=13000$, i.e., dynamic probability approaching zero around the 15th epoch, performs better than the other settings. $\omega=4600$ (dynamic probability approaching zero around the 5th epoch) results in a premature reduction in sampling probability, leading to the worst performance. Meanwhile, the performance with $\omega=17000$, while slightly inferior to that of $\omega=13000$, still surpasses the outcomes of strategies where the probability reduces to zero too prematurely. Therefore, judiciously setting the $\omega$ parameter, and learning to predict with both ground truth and previous prediction as input can effectively resolve inconsistencies between training and inference.
Figure 7. Performance on KuaiRec dataset with different hyperparameter settings in vocabulary construction.
MAE $\downarrow$ XAUC $\uparrow$
$\alpha=0.98$ $\small q_{end}=0.95$ 3.282 0.574
$\small q_{end}=0.90$ 3.216 0.601
$\small q_{end}=0.85$ 3.227 0.589
$\small q_{end}=0.80$ 3.256 0.578
$\alpha=0.99$ $\small q_{end}=0.95$ 3.253 0.590
$\small q_{end}=0.90$ 3.196 0.614
$\small q_{end}=0.85$ 3.205 0.603
$\small q_{end}=0.80$ 3.231 0.595
Figure 8. The impact of window size $n_{w}$ on MAE and XAUC for KuaiRec dataset.
A.2.4. Hyperparameter for Embedding Mixup (EM)
Here, we check the effects of the window size $w_{n}$ in Eq. (8) which control the the fixed size of window to merge representations of tokens within it, preventing significant errors. As shown in Fig. 8, we select six percentages of the vocabulary size to determine the window size values for our ablation experiments on the Kuairec dataset. The horizontal axis represents these percentages (actual window sizes), and we also expand the window size to include weighting the entire vocabulary (denoted as ”all”). The results show that as the percentage increases, performance gradually improves, reaching its peak at 50%, and further increases in the window size lead to a slight performance decline. A smaller window size may fail to merge meaningful token representations, resulting in suboptimal performance. Conversely, a larger window size may introduce noise by incorporating less relevant tokens, which can degrade performance.

### v2

@@
-Generative Regression Based Watch Time Prediction for Video Recommendation: Model and Performance
+Sequence Generation Modeling for Continuous Value Prediction
 Abstract.
-Watch time prediction (WTP) has emerged as a pivotal task in short video recommendation systems, designed to encapsulate user interests.
-Predicting users’ watch times on videos often encounters challenges, including wide value ranges and imbalanced data distributions, which can lead to significant bias when directly regressing watch time.
-Recent studies have tried to tackle these issues by converting the continuous watch time estimation into an ordinal classification task.
-While these methods are somewhat effective, they exhibit notable limitations.
-For instance, the discretization process often hinges heavily on bucket partitioning, which inherently compromises prediction accuracy.
-Moreover, the relationships existing between various partition intervals are seldom fully exploited, thereby ignoring the potential for effective error correction.
-Inspired by language modeling, we propose a novel Generative Regression (GR) paradigm for WTP based on sequence generation.
-This approach employs structural discretization to enable the lossless reconstruction of original values while maintaining prediction fidelity.
-By formulating the prediction problem as a numerical-to-sequence mapping, and with meticulously designed vocabulary and label encodings, each watch time is transformed into a sequence of tokens.
-To expedite model training, we introduce the curriculum learning with an embedding mixup strategy which can mitigate training-and-inference inconsistency associated with teacher forcing.
-We evaluate our method against state-of-the-art approaches on four public datasets and one industrial dataset.
-We also perform online A/B testing on Kuaishou, a leading video app with about 400 million DAUs, to demonstrate the real-world efficacy of our method.
-The results conclusively show that GR outperforms existing techniques significantly.
-Furthermore, we successfully apply GR to another regression task in recommendation systems, i.e., Lifetime Value (LTV) prediction, which highlights its potential as a novel and effective solution to general regression challenges.
+Continuous value prediction (CVP) plays a crucial role in short video recommendation, capturing user preferences through precise numerical estimations.
+However, traditional regression-based methods often struggle with challenges like wide value ranges and imbalanced data, leading to prediction bias.
+While ordinal classification approaches have been introduced to address these issues, their reliance on discretization reduces accuracy and overlooks inherent relationships between intervals.
+To overcome these limitations, we introduce a novel Generative Regression (GR) framework for CVP, inspired by sequence generation techniques in language modeling.
+Our method transforms numerical values into token sequences through structural discretization, preserving original data fidelity while improving prediction precision.
+Leveraging a carefully crafted vocabulary and label encoding, GR employs curriculum learning with an embedding mixup strategy to bridge training-inference gaps.
+Experimental evaluations on four public datasets and one large-scale industrial dataset validate the superiority of GR over existing methods.
+Real-world A/B tests on Kuaishou, a leading video platform, further demonstrate its practical effectiveness.
+Additionally, GR proves adaptable to other regression tasks, such as Lifetime Value (LTV) prediction, showcasing its potential as a robust solution for diverse CVP challenges.
 Recommendation, Watch-time prediction, Generative regression
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The Web Conference 2025; April 28–May 2, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems
 1. Introduction

### v3

@@
-Sequence Generation Modeling for Continuous Value Prediction
+Generative Regression Based Watch Time Prediction for Short-Video Recommendation
 Abstract.
-Continuous value prediction (CVP) plays a crucial role in short video recommendation, capturing user preferences through precise numerical estimations.
-However, traditional regression-based methods often struggle with challenges like wide value ranges and imbalanced data, leading to prediction bias.
-While ordinal classification approaches have been introduced to address these issues, their reliance on discretization reduces accuracy and overlooks inherent relationships between intervals.
-To overcome these limitations, we introduce a novel Generative Regression (GR) framework for CVP, inspired by sequence generation techniques in language modeling.
-Our method transforms numerical values into token sequences through structural discretization, preserving original data fidelity while improving prediction precision.
-Leveraging a carefully crafted vocabulary and label encoding, GR employs curriculum learning with an embedding mixup strategy to bridge training-inference gaps.
-Experimental evaluations on four public datasets and one large-scale industrial dataset validate the superiority of GR over existing methods.
-Real-world A/B tests on Kuaishou, a leading video platform, further demonstrate its practical effectiveness.
-Additionally, GR proves adaptable to other regression tasks, such as Lifetime Value (LTV) prediction, showcasing its potential as a robust solution for diverse CVP challenges.
+Watch time prediction (WTP) has emerged as a pivotal task in short video recommendation systems, designed to quantify user engagement through continuous interaction modeling.
+Predicting users’ watch times on videos often encounters fundamental challenges, including wide value ranges and imbalanced data distributions, which can lead to significant estimation bias when directly applying regression techniques.
+Recent studies have attempted to address these issues by converting the continuous watch time estimation into an ordinal regression task.
+While these methods demonstrate partial effectiveness, they exhibit notable limitations: (1) the discretization process frequently relies on bucket partitioning, inherently reducing prediction flexibility and accuracy and (2) the interdependencies among different partition intervals remain underutilized, missing opportunities for effective error correction.
+Inspired by language modeling paradigms, we propose a novel Generative Regression (GR) framework that reformulates WTP as a sequence generation task.
+Our approach employs structural discretization to enable nearly lossless value reconstruction while maintaining prediction fidelity.
+Through carefully designed vocabulary construction and label encoding schemes, each watch time is bijectively mapped to a token sequence.
+To mitigate the training-inference discrepancy caused by teacher-forcing, we introduce a curriculum learning with embedding mixup strategy that gradually transitions from guided to free-generation modes.
+We evaluate our method against state-of-the-art approaches on two public datasets and one industrial dataset.
+We also perform online A/B testing on the Kuaishou App to confirm the real-world effectiveness.
+The results conclusively show that GR outperforms existing techniques significantly.
+Furthermore, we successfully apply GR to Lifetime Value (LTV) prediction, achieving 17.66% MAE improvement over existing methods.
+These results validate GR as a generalizable solution for continuous value prediction tasks in recommendation systems.
 Recommendation, Watch-time prediction, Generative regression
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 6719ccs: Information systems Recommender systems
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
 1. Introduction
@@
 Consequently, accurate watch time estimation enables platforms to recommend videos prolonging users’ viewing, which impacts key business metrics such as Daily Active Users (DAU) and drives revenue growth.
-In contrast to limited and discrete actions such as liking, following, and sharing, watch time generally exhibits a wide range and long-tailed distribution.
-This characteristic renders it fundamentally a regression problem for prediction.
-However, the continuous nature of watch time also makes it vulnerable to outliers and potential predictive bias, thereby increasing modeling complexity and challenges of accurate prediction.
-Some methods (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a) optimize watch time prediction from a debiasing perspective but have not yet adequately addressed the core challenges of regression.
-Some others (Sun et al., 2024; Lin et al., 2023) transform the prediction problem into an Ordinal Regression (OR) task by employing a series of binary classifications across various predefined time intervals, as separately shown in Fig.
+In contrast to limited and discrete actions such as liking, following, and sharing, watch time generally exhibits a wide range and long-tailed distribution, making it fundamentally a regression problem for prediction.
+Some methods (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a; Zhang et al., 2023; Tang et al., 2023) optimize watch time prediction from a debiasing perspective but have not yet adequately addressed the core challenges of regression.
+Some others (Sun et al., 2024; Lin et al., 2023) transform the prediction problem into an Ordinal Regression (OR) task by employing a series of binary classifications across various predefined time intervals (buckets), as separately shown in Fig.
 1(a) and Fig.
 1(b).
-Nevertheless, such a modeling paradigm still exhibits two major limitations as follows:
-Firstly, conditional dependencies among time intervals are not fully leveraged.
-The dependencies are solely reflected in the definition of the labels, predictions across different time intervals are often produced independently, thereby hindering the potential for effective error correction.
-Secondly, the ordinal regression paradigm decomposes the regression target into a series of binary classification problems defined by fixed time intervals, which inherently confines the labeling of samples within these rigid divisions.
-As a result, model performance is highly contingent on the method of time interval segmentation.
-Coarse segmentation, i.e., wider spans, may overlook subtle differences, while overly fine segmentation, i.e., narrower spans, can result in insufficient data, degrading the model’s ability to learn effective features (Sun et al., 2024).
+While effective, such a modeling paradigm still exhibits two major limitations as follows:
+Firstly, conditional dependencies among time intervals are not fully leveraged, which are solely reflected in the definition of the labels.
+Predictions across different time intervals are often produced independently, thereby hindering the potential for effective error correction and leading to suboptimal results.
+We provide rigorous theoretical proof of this limitation in the supplementary material.
+Secondly, the strict discretization process within fixed time intervals in ordinal regression makes model performance highly contingent on the method of time interval segmentation, inherently reducing prediction flexibility and accuracy.
+This approach performs binary classification across all predefined buckets, with the final prediction derived as the sum of bucket sigmoid probabilities multiplied by their corresponding bucket span values.
+Due to the wide range of actual watch times (Sun et al., 2024; Tang et al., 2023), tail buckets often have excessively large span values, which can disproportionately amplify prediction errors for samples with shorter watch times, even when the binary probabilities of these tail buckets are minimal.
+Additionally, the scrolling mode of short video platforms results in a high percentage of videos with relatively short watch times in real-world scenarios, further exacerbating the overall fitting error.
 In response to these limitations above, inspired by the recent success of Large Language Models (LLMs) (Touvron et al., 2023; Brown, 2020; Zhao et al., 2023b), we propose a novel universal regression paradigm, called Generative Regression (GR), which effectively utilizes dependencies among multi-step predictions and does not strictly rely on fixed time interval divisions.
@@
 This generative regression paradigm not only ingeniously inherits the advantage of previous ordinal regression methods (Sun et al., 2024; Lin et al., 2023; Frank and Hall, 2001; Li and Lin, 2006) by decomposing the regression task into multi-classification subtasks to simplify the prediction process, but also leverages dependencies between steps to accurately and progressively approximate the total watch time.
-On the other hand, unlike ordinal regression methods (Sun et al., 2024; Lin et al., 2023) that restrict outputs to binary classification within fixed time intervals, our GR model provides the flexibility of allowing each predictive step to choose from a vocabulary of tokens, where each token represents a distinct time slot in positive real number space.
-This enlarged solution space enables GR to generate a broader set of potential sequences, thereby improving its capacity to generalize across diverse watching behaviors and leading to more accurate and personalized predictions.
-A comprehensive comparison of GR with other regression methods is shown in Tab.
-1.
-For token definition and watch time segmentation, we propose a data-driven unified vocabulary construction method, which mitigates token imbalance and eliminates reliance on manual design, and a label encoding strategy allows a lossless restoration of watch time values, thereby enhancing the model’s generality and generalization capability.
-To accelerate model convergence, we adopt curriculum learning (Bengio et al., 2009) strategy that combines teacher forcing (Venkatraman et al., 2015) with standard training to alleviate training-and-inference inconsistency, commonly known as exposure bias (Venkatraman et al., 2015; Ding and Soricut, 2017).
+On the other hand, unlike ordinal regression methods (Sun et al., 2024; Lin et al., 2023) that restrict outputs to binary classification within fixed time intervals, our GR model offers the flexibility for each predictive step to not only select from a vocabulary of tokens—each representing a distinct time slot in positive real number space, but also output an end-of-sequence (¡EOS¿) token.
+This flexibility enables GR to generate a broader set of potential sequences, thereby improving its capacity to generalize across diverse watching behaviors and leading to more accurate and personalized predictions.
+For token definition and watch time segmentation, we propose a data-driven unified vocabulary construction method, which mitigates token imbalance and eliminates manual design reliance, and a label encoding strategy allows a lossless restoration of watch time values, thereby enhancing the model’s generality and generalization capability.
+To accelerate model convergence, we adopt curriculum learning (Bengio et al., 2009) strategy during training to alleviate training-and-inference inconsistency, commonly known as exposure bias (Venkatraman et al., 2015; Ding and Soricut, 2017).
 Besides, leveraging our insights into the training process, we propose an embedding mixup method to compensate for output-to-input gradients.
 This approach enhances model performance at a lower computational cost by leveraging the semantic additivity of tokens while ensuring consistency between training and inference.
-Table 1.
-A comprehensive comparison with existing deep regression methods.
-‘H’, ‘M’, and ‘L’ indicate high, medium, and low respectively.
-Complexity represents computation overhead during inference, with $m$, $n$, $d$, and $t$ indicating the last hidden feature size, discretization segments, vocabulary size, and sequence length, respectively.
-Methods Predictive Paradigm Discre -tization Accuracy Versatility Complexity
-VR Discriminative - L H $\mathcal{O}(m)$
-TPM (Lin et al., 2023) Discriminative Tree-based M M $\mathcal{O}(mn)$
-CREAD (Sun et al., 2024) Discriminative Linear M M $\mathcal{O}(mn)$
-GR Generative Sequential H H $\mathcal{O}(mdt)$
 The contributions of this paper are as follows:
@@
   2. (ii)
-    To enhance the generality and adaptability of GR, we develop a data-driven unified vocabulary design and a label encoding method.
-Additionally, we introduce curriculum learning with embedding mixup to mitigate exposure bias and accelerate model training.
+    To enhance generality and adaptability, we develop a data-driven unified vocabulary design and a label encoding method.
+Additionally, we introduce curriculum learning with embedding mixup to mitigate exposure bias and compensate for output-to-input gradients to accelerate model training.
   3. (iii)
-    Extensive online and offline experiments demonstrate that GR significantly outperforms existing SOTA models.
-We further analyze the impact of key factors like vocabulary design and model architecture, to provide a clear understanding of the mechanisms and processes underlying GR.
+    Extensive online and offline experiments show that GR significantly outperforms existing SOTA models.
+We further analyze the underlying reasons for performance gain and the impact of key factors like vocabulary design to provide a clear understanding of the mechanisms underlying GR.
   4. (iv)
@@
 2. Related Work
-Watch time prediction aims to estimate the video watch time based on the user’s profile, historical interactions, and video characteristics.
+2.1. Watch Time Prediction (WTP)
+WTP aims to estimate the video watch time based on the user’s profile, historical interactions, and video characteristics.
 Value regression (VR) directly predicts the absolute value of watch time, assessing model accuracy by mean square error (MSE).
-WLR (Covington et al., 2016) applies watch duration as weights within its loss function, yet it may suffer from significant bias due to its weighting mechanism.
-D2Q (Zhan et al., 2022) segments data by duration, predicting watch time quantiles within these segments to alleviate duration bias.
+Subsequent WTP methods can be roughly divided into two groups.
+The first focuses on optimizing WTP from a debiasing perspective (Zhan et al., 2022; Zhao et al., 2024; Zheng et al., 2022; Zhao et al., 2023a; Zhang et al., 2023; Tang et al., 2023).
 CWM (Zhao et al., 2024) introduces a counterfactual watch time, estimating a video’s hypothetical full watch time to gauge user interest.
@@
 However, these methods have not yet adequately addressed the core challenges of regression.
-Some solutions transform the problem into ordinal regression.
+The second transforms the regression task into classification (Sun et al., 2024; Lin et al., 2023; Covington et al., 2016).
 CREAD (Sun et al., 2024) introduces an error-adaptive discretization technique to construct dynamic time intervals.
@@
 Yet, these approaches are unable to fully capitalize on the interdependencies among these intervals and heavily rely on time interval segmentation.
-Ordinal regression, also known as ordinal classification, is a type of predictive modeling strategy employed when the outcome variable is ordinal and the relative order of labels is important, such as age prediction (Niu et al., 2016), monocular depth perception (Fu et al., 2018), and head-pose estimation (Hsu et al., 2018).
-CNNOR (Liu et al., 2017) introduces a deep learning model for ordinal regression tailored to small datasets.
-SORD (Diaz and Marathe, 2019) employs a soft ordinal label approach to train the ordinal regression model.
-Additionally, (Li et al., 2021) crafts a probabilistic embedding strategy specifically for ordinal regression tasks.
-Ordinal regression has not been applied to watch time prediction until the introduction of CREAD (Sun et al., 2024) and TPM (Lin et al., 2023).
-These models follow the traditional approach of decomposing regression into multiple binary classification tasks, achieving significant benefits.
+2.2. Ordinal Regression
+OR is a type of predictive modeling strategy employed when the outcome variable is ordinal and the relative order of labels is important, such as age prediction (Niu et al., 2016), monocular depth perception (Fu et al., 2018), and head-pose estimation (Hsu et al., 2018).
+Recent works include specialized architectures like CNNOR (Liu et al., 2017), alternative training paradigms using soft labels such as SORD (Diaz and Marathe, 2019), and dedicated probabilistic embedding methods (Li et al., 2021).
+It has not been applied to watch time prediction until the introduction of CREAD (Sun et al., 2024) and TPM (Lin et al., 2023).
+These models decompose the regression task into multiple binary classification tasks, achieving significant benefits.
 Figure 2.
-The framework of GR model, which adopts an encoder-decoder RNN-based architecture.
-The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner and employs the curriculum learning with embedding mixup (CLME) strategy to alleviate training-and-inference inconsistency introduced by teacher forcing.
-Sequence generation requires the model to understand the context of an input sequence and generate an appropriate output sequence.
-Early researches focus on natural language processing (NLP) tasks such as machine translation (Sutskever, 2014; Elman, 1990; Graves and Graves, 2012; Cho, 2014), text summarization (Bahdanau, 2014; Vaswani, 2017), and dialogue generation (Devlin, 2018; Raffel et al., 2020).
+The framework of the GR model, which adopts an encoder-decoder architecture.
+The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner and employs the curriculum learning with embedding mixup (CLEM) strategy to alleviate training-and-inference inconsistency introduced by teacher forcing.
+2.3. Sequence Generation
+Sequence generation learns contextual sequence mappings, initially prominent in NLP for tasks like machine translation (Sutskever, 2014; Cho, 2014) and text summarization (Bahdanau, 2014; Vaswani et al., 2017).
+This paradigm extended to recommendation systems for capturing sequential user behavior patterns.
 In recommendation systems, sequential recommendation methods have been proposed to capture sequential patterns.
@@
 BERT4Rec (Sun et al., 2019) employs a bidirectional transformer to encode item sequences.
-However, sequence generation models in recommendation systems predominantly focus on predicting the sequence of user behaviors, and their application to watch time prediction remains unexplored.
+However, these sequential recommendation methods have predominantly focused on predicting the sequence of user behaviors, and their application to watch time prediction remains unexplored.
 3. Method
@@
 tags, duration and category) of the $i$-th example respectively   [^1] We omit the context-side features for simplicity., $y_{i}\in\mathbb{R}$ is the corresponding watch time of the $i$-th example collected from recommendation system logs.
-Value regression (VR) methods aim to learn a function $f(\cdot)$ that directly maps the input features to a real-valued output, i.e., $y_{i}=f([\bm{u_{i}};\bm{v_{i}}])$.
+Value regression methods aim to learn a function $f(\cdot)$ that directly maps the input features to a real-valued output, i.e., $y_{i}=f([\bm{u_{i}};\bm{v_{i}}])$.
 Sequence generation in GR is mostly based on autoregressive language modeling.
-Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$, where each element $w_{j}$ represents a predefined time slot (e.g.
+Specifically, we introduce a vocabulary $\mathcal{V}=\left\{w_{j}\right\}_{j=1}^{V}$, where $V$ is the vocabulary size and each element $w_{j}$ represents a predefined time slot (e.g.
 5 seconds, 10 seconds, etc.).
@@
 Thus, “token” and “time slot” will be used interchangeably in the sequel.
-The vocabulary embedding matrix is denoted as $\bm{E}\in\mathbb{R}^{V\times D}$, where $V$ is the vocabulary size, and $D$ is the dimension of the time slot embeddings.
-We decompose $y_{i}$ into a sequence of tokens $\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}$, where $s_{i}^{t}\in\mathcal{V}$ and $T_{i}$ is the length of the sequence.
-We call this process label encoding, which is detailed in Sec.
+The vocabulary embedding matrix is denoted as $\bm{E}\in\mathbb{R}^{V\times D}$, where $D$ is the dimension of the time slot embeddings.
+We decompose $y_{i}$ into a sequence of tokens $\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}$, where $s_{i}^{t}\in\mathcal{V}$ and $T_{i}$ denotes the length of the sequence.
+This process, referred to as label encoding, is described in detail in Sec.
 3.4.
-On the other hand, we design a label decoding function $\text{g}(\cdot)$ such that derives the original watch time $y_{i}$ from $\bm{s_{i}}$, i.e., $y_{i}=\text{g}(\bm{s_{i}})=\sum_{t=1}^{T_{i}}\text{g}(s_{i}^{t})\in\mathbb{R}$.
-Our goal is to train a sequence generation model, given the user and video features $(\bm{u_{i}},\bm{v_{i}})$, generates the corresponding sequence of watch time slots $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, and in turn, from which the predicted watch time $\hat{y_{i}}=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$ approximates the actual watch time $y_{i}$.
+On the other hand, we design a label decoding function $\text{g}(\cdot)$  [^2] Here, $\text{g}(\cdot)$ functions as a lookup table that maps tokens to real-valued vocabulary entries, e.g., $\text{g}(``\text{30s}")=30$.
+that reconstructs the original watch time $y_{i}$ from $\bm{s_{i}}$, i.e., $y_{i}=\text{g}(\bm{s_{i}})=\sum_{t=1}^{T_{i}}\text{g}(s_{i}^{t})\in\mathbb{R}$.
+Our goal is to train a sequence generation model, given user and video characteristics $(\bm{u_{i}},\bm{v_{i}})$, which generates the corresponding sequence of watch time slots $\bm{\hat{s}_{i}}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, and in turn, from which the predicted watch time $\hat{y_{i}}=\text{g}(\bm{\hat{s}_{i}})=\sum_{t=1}^{T_{i}}\text{g}(\hat{s}_{i}^{t})$ approximates the actual watch time $y_{i}$.
 3.2. The Generative Regression (GR) Model
-As shown in Fig. 2, GR adopts an encoder-decoder RNN-based architecture. The encoder extracts user and video features, while the decoder predicts watch time in an autoregressive manner.
+As shown in Fig. 2, GR adopts a Transformer-based encoder-decoder architecture. The encoder extracts user and video features, while the decoder predicts the watch time in an autoregressive manner.
 3.2.1. Encoder
-Unlike traditional sequence-to-sequence tasks or user behavior modeling in recommendation systems, watch time prediction does not inherently depend on history action sequences.
-Consequently, directly using an RNN as the encoder — designed for processing temporally dependent sequential data — is inappropriate and may result in suboptimal feature representation.
-To ensure the generality and simplicity of the model, we use a feed-forward network (FFN) as the encoder.
-However, it can be replaced with any complex model architecture.
+Unlike traditional sequence-to-sequence tasks or user behavior modeling in recommendation systems, watch time prediction does not inherently depend on the order of user history interacted items.
+To ensure model generality and simplicity, we follow previous works (Lin et al., 2023; Sun et al., 2024) and employ a feedforward network (FFN) as an encoder.
+Note that this encoder can be replaced with any sophisticated model architecture.
 Formally, the encoder extracts user and video features to produce a fixed-length hidden feature $\bm{h_{i}}\in\mathbb{R}^{1\times D}$ that will be fed to the decoder as follows:
@@
 (2) $\hat{s}_{i}^{t}=\arg\max_{w\in\mathcal{V}}P_{\theta}(w\mid\bm{h_{i}},\bm{\hat{s}_{i}^{<t}})$
-where $\theta$ is the model parameter and $\bm{\hat{s}_{i}^{<t}}$ is the tokens generated before. Utilizing the chain rule, the overall probability of generating the sequence can be expressed as
-(3) $P_{\theta}(\bm{s_{i}}\mid\bm{h_{i}})=P_{\theta}(s_{i}^{1},...,s_{i}^{T_{i}}\mid\bm{h_{i}})=\prod_{t=1}^{T}P(s_{i}^{t}\mid\bm{h_{i}},\bm{{s}_{i}^{<t}})$
-There remain three issues to be addressed: how to construct an effective vocabulary?
-how to encode $y_{i}$ into a sequence $\bm{s_{i}}$?
-and how to optimize the model?
-which are detailed in the following sections.
+where $\theta$ is the model parameter and $\bm{\hat{s}_{i}^{<t}}$ represents the tokens generated before.
+Utilizing the chain rule, the overall probability of generating the sequence can be expressed as
+(3) $P_{\theta}(\bm{s_{i}}\mid\bm{h_{i}})=P_{\theta}(s_{i}^{1},...,s_{i}^{T_{i}}\mid\bm{h_{i}})=\prod_{t=1}^{T}P_{\theta}(s_{i}^{t}\mid\bm{h_{i}},\bm{{s}_{i}^{<t}})$
+Three key issues remain to be addressed: (1) how to construct an effective vocabulary, (2) how to encode $y_{i}$ into a sequence $\bm{s_{i}}$, and (3) how to optimize the model.
+These issues are detailed in the following sections.
 3.3. Vocabulary Construction
@@
 Based on our cognition of the deep regression task, three principles are designed to guide the construction of vocabulary.
-  • Completeness: The vocabulary $\mathcal{V}$ must be able to represent all watch time values $\{y_{i}\}_{i=1}^{N}$ using a finite number of tokens almost without loss [^2] The label space is continuous, it is impossible to recover the original watch time from a discretized sequence with zero error..
-Also, each token must be unique.
+  • Completeness: The vocabulary $\mathcal{V}$ must be able to represent all watch time values $\{y_{i}\}_{i=1}^{N}$ using a finite number of tokens almost without loss.Also, each token must be unique.
   • Balance: The frequencies of tokens should be relatively uniform to prevent class imbalance.
@@
 An alternative is to select watch time values as tokens based on one fixed percentile, then subtract the token values from all watch time values that exceed them, repeating this process until the residuals become negligible, which fails to meet the balance principle.
+Due to the space limit, details of this strategy are provided in the supplementary materials.
 Algorithm 1 Constructing Vocabulary with dynamic percentiles
-1:Dataset labels $\bm{Y}=\{y_{j}\}_{j=1}^{N}$, initially empty Vocabulary $\mathcal{V}=\{\}$, precision threshold $\epsilon$, start percentile $q_{\text{start}}$, end percentile $q_{\text{end}}$, decay rate $\alpha$, minimal restoration error $\epsilon$.
-2:Sort $\bm{Y}$ in descending order to obtain $\hat{\bm{Y}}$.
+1:Dataset labels $\bm{Y}=\{y_{j}\}_{j=1}^{N}$, initially empty Vocabulary $\mathcal{V}=\{\}$, start percentile $q_{\text{start}}$, end percentile $q_{\text{end}}$, decay rate $\alpha$, minimal restoration error $\epsilon$.
+2:Sort $\bm{Y}$ in descending order to obtain $\hat{\bm{Y}}=\{\hat{y}_{j}\}_{j=1}^{N}$.
 3:Initialize iteration counter $i=1$, error metric $err=\infty$, current percentile $q=q_{\text{start}}$
 4:while $err>\epsilon$ do
-5:     Compute the $q$-percentile $z_{i}$ of $\bm{Y}$.
-6:     if $z_{i}=0$ then $\triangleright$ Terminate if the percentile value is zero
+5:    Compute the $q$-percentile $o_{i}$ of $\hat{\bm{Y}}$.
+6:    if $o_{i}=0$ then $\triangleright$ Terminate if the percentile value is zero
 7:         break
 8:    end if
-9:     Insert $z_{i}$ into vocabulary $\mathcal{V}$.
+9:    Generate a new token $v_{i}$ which satisfy $o_{i}=g(v_{i})$ and insert $v_{i}$ into vocabulary $\mathcal{V}$.
 10:    Update $\hat{\bm{Y}}$ using:
-$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if }\hat{y}_{j}<z_{i},\\ \hat{y}_{j}-z_{i},&\text{otherwise}\end{cases}$
+$\hat{y}_{j}=\begin{cases}\hat{y}_{j},&\text{if }\hat{y}_{j}<o_{i},\\ \hat{y}_{j}-o_{i},&\text{otherwise}\end{cases}$
 11:    Update the error metric $err$: $err=\max\{\frac{\hat{y}_{j}}{y_{j}}\}_{j=1}^{N}$
@@
 15:return $\mathcal{V}$
-To meet these principles above, we develop a data-drived vocabulary construction algorithm based on dynamic percentile, detailed in Algorithm.
-1.
-Specifically, the algorithm starts with an initially large percentile $q_{start}$ and decreases it according to the decay rate $\alpha$, ensuring that the percentile does not fall below the endpoint $q_{end}$.
-This strategy expedites the reduction of tail values, rapidly decreasing the variance among updated values, which effectively mitigates the challenges posed by the long-tailed distribution in the dataset, for which we provide detailed validation in the experimental section.
-Under the generative regression framework, it is important to note that the process of constructing a vocabulary and converting the task target (e.g.
-watch time) into a sequential token format (a.k.a label encoding) resembles the syntax construction of a language.
-However, we do not contend that the aforementioned strategy is the most optimal one.
+To address both principles simultaneously, we propose a data-driven vocabulary construction algorithm using dynamic quantile adjustment (Algorithm.
+1).
+The algorithm initializes with a high starting quantile $q_{start}$ and adaptively reduces it by decay rate $\alpha$ until reaching the terminal quantile $q_{end}$.
+This strategy expedites the reduction of tail values, rapidly decreasing the variance among updated values, which effectively mitigates the challenges posed by the long-tailed distribution in the dataset, for which we provide detailed experimental validation in Sec.
+4.4.
+We emphasize that our vocabulary construction and label encoding process, while analogous to linguistic syntax building for sequence generation, does not presume theoretical optimality.
+The proposed strategy serves as a principled engineering solution, leaving theoretical analysis of optimal tokenization for future work.
 3.4. Label Encoding
-Given the vocabulary $\mathcal{V}$, we need to do label encoding, i.e., transforming the watch time values $\{y_{i}\}_{i=1}^{N}$ to corresponding target sequences $\{\bm{s_{i}}=\{s_{i}^{1},...,s_{i}^{T_{i}}\}\}_{i=1}^{N}$.
-To facilitate the learning process with the encoded labels (i.e., transformed token sequences), we also offer three principles for label encoding as follows:
-  • Correctness: $y_{i}=\sum_{t=0}^{T_{i}}w_{i}^{t}+\epsilon$, where $\epsilon$ is a small value (e.g. 0.1% relative error) which can be omitted.
-  • Minimal sequence length: $T_{i}$ is the possible minimal length of the token sequence selected to satisfy correctness for $y_{i}$.
-  • Monotonicity: $s_{i}^{1}\geq s_{i}^{2}\geq...\geq s_{i}^{T_{i}}$ is required.
-The minimum sequence length requirement simplifies the learning process.
-Additionally, the non-decreasing order captures the dynamic changes in user interest while watching videos, making the sequence generation process progressively (from easy to hard).
-Without loss of generality, the token indices in the vocabulary are also sorted in descending numerical order.
-To this end, we employ a greedy strategy for the shortest sequence length, starting from the largest possible watch time slot and decreasing progressively, decomposing the total watch time $y_{i}$ into a sequence of watch time slots.
+Given the vocabulary $\mathcal{V}=\{w_{1},w_{2},...,w_{V}\}$, we perform label encoding to transform the watch time values $\{y_{i}\}_{i=1}^{N}$ into corresponding target sequences $\{\bm{s}_{i}=\{s_{i}^{1},\ldots,s_{i}^{T_{i}}\}\}_{i=1}^{N}$.
+To guide the label encoding process, we propose three foundational principles:
+  • Correctness: The original value must be reconstructible from the token sequence with bounded error:
+    (4) $y_{i}=\sum_{t=1}^{T_{i}}g(s_{i}^{t})+\epsilon,\quad\text{where }|\epsilon|\leq 0.001\cdot y_{i}$
+  • Minimal Sequence Length: The sequence length $T_{i}$ should achieve the minimal possible cardinality while satisfying the correctness constraint.
+  • Monotonicity: Token values must satisfy a non-increasing order:
+    (5) $g(s_{i}^{1})\geq g(s_{i}^{2})\geq\cdots\geq g(s_{i}^{T_{i}})$
+The minimum sequence length principle reduces learning complexity, while the monotonic constraint captures decaying user attention patterns during video watching.
+To follow these principles, we implement a greedy decomposition algorithm.
+Starting from the largest possible watch time slot and decreasing progressively, decomposing the total watch time $y_{i}$ into a sequence of watch time slots.
 3.5. Optimization and Inference
 3.5.1. Vanilla Training Process
-Similar to language modeling, the model predicts the next token $s^{t}$ using the ground truth tokens $s^{<t}$.
-The learning objective is to maximize the negative log-likelihood of the generated sequence over the training set and effectively minimize the difference between the predicted sequence and the ground truth sequence, which is computed by the cross-entropy loss as follows:
+Following language modeling paradigms, the model predicts the next token $s^{t}$ conditioned on preceding ground truth tokens $s^{<t}$.
+The learning objective minimizes the cross-entropy loss between predicted and ground truth sequences:
 (6) $\mathcal{L}_{ce}=-\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log P_{\theta}(\hat{s}_{i}^{t}\mid\bm{h_{i}},\hat{s}_{i}^{<t})$
@@
 where $\hat{y_{i}}=\sum_{t=1}^{T_{i}}g(\hat{s}_{i}^{t})$, $\delta$ acts as a threshold, toggling between quadratic and linear losses to balance sensitivity and robustness against outliers.
-Therefore, the loss function we need to optimize is
+Therefore, the composite loss becomes:
 (8) $\mathcal{L}=\mathcal{L}_{ce}+\lambda\cdot\mathcal{L}_{huber}$
-$\lambda$ is a hyperparameter that balances the two losses.
-To improve model efficiency, we adopt a teacher forcing (TF) strategy, which directly feeds the ground truth output $s_{i}^{t}$ as input at step $t+1$ to guide model training.
-However, since the ground truth is unknown during inference, the discrepancy in input for the decoder leads to the well-known exposure bias problem, which can degrade model performance (Goodman et al., 2020).
+where $\lambda$ is a hyperparameter that balances the two losses.
+To improve model efficiency, we adopt a teacher forcing (TF) strategy (Venkatraman et al., 2015), which directly feeds the ground truth output $s_{i}^{t}$ as input at step $t+1$ to guide model training.
+However, since the ground truth is unknown during inference, the discrepancy of input for the decoder leads to the well-known exposure bias problem (Goodman et al., 2020), which can degrade model performance.
 3.5.2. Curriculum Learning with Embedding Mixup (CLEM)
-We introduce a Curriculum Learning (CL) strategy that combines TF and vanilla training (a.k.a free running mode) by progressively incorporating stochasticity into the training process.
-Specifically, to predict $\hat{s}_{i}^{t}$, we randomly choose ground truth token $s_{i}^{t-1}$ with a dynamic probability $p$ as the sampling rate.
+To mitigate exposure bias inherent in teacher forcing, we propose a phased Curriculum Learning (CL) strategy.
+Specifically, to predict $\hat{s}_{i}^{t}$, we randomly choose ground truth tokens $s_{i}^{t-1}$ or predicted tokens $\hat{s}_{i}^{t-1}$ with a dynamic probability $p$ as the sampling rate.
+However, Transformer processes the entire sequence in parallel during a single forward pass, preventing access to the predicted tokens of previous time steps.
+Thus, as shown in Fig.
+2, we implement CL with two forward passes through the decoder during training.
+The first pass performs vanilla training to obtain initial model predictions.
+In the second pass, inputs are sampled between ground truth tokens and predicted tokens with probability $p$, yielding the final predictions.
+Both passes share the same model parameters.
 To warm up, we start with $p\approx 1$, indicating that the model predominantly relies on the ground truth tokens.
@@
 Figure 3. Watch time embedding with a weighted sum of token embeddings (left) and the probability distribution difference among tokens for each $\hat{s}_{i}^{t}$ (right). Best viewed in color.
-We notice that GR can capture the inherent relationship between different tokens.
-As the cardinality of $\mathcal{V}$ is not as large as that in LMs, it is hard to visualize $\mathcal{V}$’s embedding matrix $\bm{E}$.
-As an alternative, we visualize the target numerical embeddings of $y_{i}$ by $\bm{e_{i}}=\sum_{t=1}^{T_{i}}r_{t}\bm{E}[\hat{s_{i}}^{t},:]$, $r_{t}=\frac{g(\hat{s}_{i}^{t})}{y_{i}}$ is the contribution of $\hat{s}_{i}^{t}$ to the target value $y_{i}$.
-As shown in Fig.
-3(a), values beginning with the same token $w_{j}$ are clustered, and the initial tokens of similar numeric values $\text{g}(w_{j})$ are also relatively close to each other.
-This clustering feature indicates a coherent structure in the embedding space, suggesting that the model effectively captures similarities among the tokens.
-Although this is what we expect the model to learn, it can also lead to the issue of confusion among similar tokens.
-In other words, the model may struggle to effectively distinguish between the target token and its neighboring tokens.
+Our analysis reveals that GR effectively captures inter-token relationships through its embedding structure.
+Given the vocabulary size $V$ being orders of magnitude smaller than typical language models, we analyze token semantics via aggregated value embeddings:
+(10) $\bm{e}_{i}=\sum_{t=1}^{T_{i}}r_{t}\bm{E}[s_{i}^{t},:],\quad r_{t}=\frac{g(s_{i}^{t})}{y_{i}}$
+where $r_{t}$ represents the contribution of token $s_{i}^{t}$ to target value $y_{i}$. Fig. 3(a) demonstrates two key properties:
+  • Token Clustering: Values sharing initial tokens form distinct clusters.
+  • Semantic Continuity: Embeddings of tokens with similar $g(w_{j})$ values reside in proximate regions.
+This structural coherence facilitates numerical reasoning through geometrically meaningful representations.
 As noted in Sec.
 3.4, tokens are arranged in non-increasing order within the vocabulary $\text{g}(w_{1})>\text{g}(w_{2})>...>\text{g}(w_{V})$.
-We illustrate this phenomenon in Fig.
-3(b) by computing the averaged probability difference of each token relative to its neighbors.
-The results indicate that the predicted probability of the target token is closely aligned with the surrounding tokens, which contributes to an increase in the final prediction error.
-Since our token sequences are encoded in non-increasing order, it becomes challenging for subsequent tokens to bridge the gap between $\hat{y}_{i}$ and ground truth $y_{i}$ once the first token, denoted as $\hat{s_{i}}^{1}$ is wrong.
-Additionally, the argmax operation in Eq.
-(2) will disrupt the gradient propagation path, which further complicates the learning process.
-We observe that employing a curriculum training strategy with inference by using a beam search ensemble significantly boosts model performance.
-However, it incurs more computational cost.
-Given the previously discussed properties of tokens and sequences, the numerical differences between adjacent tokens are minimal.
-To address this, we propose to integrate the embedding sequences of the preceding tokens through an ensemble approach during the training process to improve the prediction of the next token.
-Embedding Mixup (EM).
-The cohort is centered on the current predicted token $\hat{s}^{t}_{i}$ with window size $n_{w}$, the mixup region is $[\hat{s}^{t}_{i}-b,\hat{s}^{t}_{i}+b]$ and $b=\lfloor\frac{n_{w}}{2}\rfloor$.
+We also compute the averaged probability difference of each token relative to its neighbors and observe that tokens with neighboring indices in the vocabulary demonstrate the highest probability similarity in the model’s predictions, as shown in Fig.
+3(b).
+To improve the prediction precision of next token, we propose to integrate the embedding sequences of the preceding tokens through a local ensemble approach called Embedding Mixup (EM) during the training process.
+The cohort is centered on the current predicted token $\hat{s}^{t}_{i}$ with window size $n_{w}$, the mixup region is $[\delta_{\hat{s}^{t}_{i}}-b,\delta_{\hat{s}^{t}_{i}}+b]$ and $b=\lfloor\frac{n_{w}}{2}\rfloor$, $\delta_{\hat{s}^{t}_{i}}$ represents the token index of $\hat{s}^{t}_{i}$ in $\mathcal{V}$.
 We have
-(8) $\bm{z_{i}}^{t}=\sum\limits_{j=0}^{n_{w}}\sigma_{j}\cdot\bm{E}[\hat{s}^{t}_{i}+j-b,:]$
-(9) $\sigma_{j}=\frac{exp(-\rho_{j})}{\sum_{k=0}^{n_{w}}exp(\rho_{\hat{s}^{t}_{i}+k-b})}$
-where $\sigma_{j}$ recalculates the fusion weights of tokens in the fixed window size, $\rho_{j}$ is the logit predicted by the decoder at step $t$.
-And $\bm{z_{i}}^{t}$ will replace original $\bm{E}[\hat{s}^{t}_{i},:]$ as the input for the next step.
-This re-weighted mixup approach leverages the numerical semantics and additivity property inherent in our tokens.
+(11) $\bm{z_{i}}^{t}=\sum\limits_{j=0}^{n_{w}}\sigma_{j}\cdot\bm{E}[\delta_{\hat{s}^{t}_{i}}+j-b,:]$
+(12) $\sigma_{j}=\frac{exp(-\rho_{j})}{\sum_{k=0}^{n_{w}}exp(\rho_{\delta_{\hat{s}^{t}_{i}}+k-b})}$
+where $\bm{E}\in\mathbb{R}^{V\times D}$ is the vocabulary embedding matrix, $\sigma_{j}$ recalculates the fusion weights of tokens in the fixed window size, $\rho_{j}$ is the logit predicted by the decoder at step $t$.
+Specifically, $\bm{z_{i}}^{t}$ will replace original $\bm{E}[\hat{s}^{t}_{i},:]$ as the input at $t+1$ step.
+This re-weighted EM approach leverages the numerical semantics and additivity property inherent in our tokens.
 The re-weighting ensures that the semantic space is aligned, eliminating any discrepancies in scale.
-EM offers three benefits: 1) It reduces the learning complexity of the model by merging representations of tokens within a fixed window, thereby preventing significant errors; 2) The integration leverages the predicted scores from the previous steps, enhancing the information transfer from output to input in recurrence structure; and 3) it ensures consistency between training and inference while lowering inference cost.
+EM offers three benefits: 1) It reduces the learning complexity of the model by merging representations of tokens within a fixed window, thereby preventing significant errors; 2) The integration leverages the predicted scores from the previous steps, enhancing the information transfer from output to input in recurrence structure and restructuring the gradient propagation path; and 3) it ensures consistency between training and inference while lowering inference cost.
 3.5.3. Inference Process
-During inference, the encoder extracts $\bm{h_{i}}$ from input features $[\bm{u_{i}},\bm{v_{i}}]$, the decoder begins with the ¡SOS¿ token, and sequentially generates the prediction sequence $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$.
+During inference, the encoder extracts $\bm{h_{i}}$ from input features $[\bm{u_{i}},\bm{v_{i}}]$, the decoder begins with the ¡SOS¿ token and sequentially generates the prediction sequence $\hat{s}_{i}=\{\hat{s}_{i}^{1},\hat{s}_{i}^{2},...,\hat{s}_{i}^{T_{i}}\}$, with each token generated using only the first forward pass.
 The process continues until the token ¡EOS¿ is generated, which signifies the completion of the sequence.
@@
   • RQ1: How does GR compare to state-of-the-art methods in terms of prediction accuracy of watch time?
-  • RQ2: What is the effect of vocabulary design on the performance of GR and why?
-  • RQ3: How do various neural network architectures perform within the GR framework?
+  • RQ2: What are the underlying reasons behind the model’s performance exceeding the baseline?
+  • RQ3: What is the effect of vocabulary design on the performance of GR and why?
   • RQ4: What impact does CLEM have on the GR model, and how do different training strategies affect performance?
   • RQ5: How does GR perform on other regression tasks?
-Due to the space limit, additional ablation study results are available at A.2.
 4.1. Experiment Settings
-4.1.1. Datasets
-Two publicly available datasets and one industrial dataset are used to evaluate the proposed method.
-The large-scale industrial dataset (Indust for short) is sourced from a real-world streaming short-video app Kuaishou with about 400 million DAUs and multi-billion impressions each day.
-We collect interaction logs for 4 days and utilize the subsequent day’s data for evaluation.
-The CIKM16333https://competitions.codalab.org/competitions/11161, sourced from the CIKM16 Cup competition, is designed to predict user engagement duration in online search sessions.
-It contains 310,302 sessions, 122,991 items, and an average session length of 3.981.
-KuaiRec (Gao et al., 2022a) is a real-world dataset collected from the video view logs of the Kuaishou app, comprising data from 7,176 users, 10,728 items, and 12,530,806 impressions.
-In addition to directly predicting watch time, some researchers (Lin et al., 2023) focus on predicting the watch ratio, which can be used in conjunction with video duration to calculate watch time.
-Therefore, we report both watch time and watch ratio results on the KuaiRec dataset.
-Table 2. Performance comparison among different approaches on KuaiRec, CIKM16 and Indust.
+4.1.1. Datasets.
+We evaluate our method on one industrial dataset and two public benchmarks.
+The large-scale industrial dataset (Indust for short) is sourced from a real-world short-video app Kuaishou with over 400 million DAUs and multi-billion impressions each day.
+We use interaction logs spanning 4 days for training and those from the subsequent day for testing.
+We also use the public CIKM16333https://competitions.codalab.org/competitions/11161 and KuaiRec (Gao et al., 2022a) datasets, adopting the experimental settings from previous works (Lin et al., 2023; Sun et al., 2024) (Details are provided in the supplementary material).
+Consistent with prior work (Lin et al., 2023), we also report watch ratio results on KuaiRec, which can be used in conjunction with video duration to calculate watch time.
+4.1.2. Metrics
+To evaluate the proposed method, we follow previous work (Lin et al., 2023; Sun et al., 2024) and utilize two performance metrics:
+  • Mean Average Error (MAE): It quantifies regression accuracy by averaging the absolute deviations between predicted values $\{\hat{y_{i}}\}_{i=1}^{N}$ and actual values $\{y_{i}\}_{i=1}^{N}$ by $\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$.
+  • XAUC (Zhan et al., 2022): This measure assesses the concordance between the predicted and actual ordering of watch time values.
+We uniformly sample pairs from the test set and calculate the XAUC by determining the percentile of samples that are correctly ordered.
+A higher XAUC indicates better model performance.
+Table 1. Performance comparison among different approaches on KuaiRec, CIKM16 and Indust dataset.
 Method KuaiRec (watch time) KuaiRec (watch ratio) CIKM16 Indust
 MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$ XAUC Improv. MAE $\downarrow$ XAUC $\uparrow$
-VR 7.634 0.504 - 0.385 0.691 - 1.039 0.641 - 46.343 0.588
-WLR (Covington et al., 2016) 6.047 0.525 4.167% 0.375 0.698 1.013% 0.998 0.672 4.836% - -
-D2Q (Zhan et al., 2022) 5.426 0.565 12.103% 0.371 0.712 3.039% 0.899 0.661 3.120% - -
-TPM (Lin et al., 2023) 3.456 0.571 13.294% 0.361 0.734 6.223% 0.850 0.676 5.460% 41.486 0.593
-CREAD (Sun et al., 2024) 3.307 0.594 17.857% 0.369 0.738 6.802% 0.865 0.678 5.772% 39.979 0.597
-GR (ours) 3.196 0.614 21.825% 0.333 0.753 8.972% 0.815 0.691 7.80% 38.528 0.604
+VR 7.634 0.534 - 0.385 0.691 - 1.039 0.641 - 46.343 0.588
+WLR (Covington et al., 2016) 6.047 0.545 2.059% 0.375 0.698 1.013% 0.998 0.672 4.836% - -
+D2Q (Zhan et al., 2022) 5.426 0.565 8.757% 0.371 0.712 3.039% 0.899 0.661 3.120% - -
+CWM (Zhao et al., 2024) 3.452 0.580 8.614% 0.368 0.725 4.920% 0.891 0.662 3.276 % - -
+TPM (Lin et al., 2023) 3.456 0.571 6.929% 0.361 0.734 6.223% 0.850 0.676 5.460% 41.486 0.593
+CREAD (Sun et al., 2024) 3.307 0.594 11.236% 0.369 0.738 6.802% 0.865 0.678 5.772% 39.979 0.597
+GR (ours) 3.196 0.614 14.981% 0.333 0.753 8.972% 0.815 0.691 7.80% 38.528 0.604
     Here, the best and second best results are marked in bold and underline, respectively.
@@
 Table 2. Performance gain on online A/B testing.
-A/B test APP usage time Average app usage per user
-+0.055% +0.072%
-APP launch counts Video consumption time
-+0.042% +0.129%
+A/B test APP Usage Time +0.112% (p-value=0.01)
+Average App Usage Per User +0.087%
+Video Consumption Time +0.129%
+    In a stable video recommendation system, a 0.1% increase is significant.
+Table 3. Comparison of vocabulary construction methods.
+Vocabulary design KuaiRec CIKM16
+MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
+Manual 3.281 0.604 0.825 0.685
+Binary 3.268 0.605 0.821 0.687
+Dynamic quantile 3.196 0.614 0.815 0.691
 4.1.3. Compared Methods
-Considering baseline methods compared in prior studies (Sun et al., 2024; Lin et al., 2023), we compare several state-of-the-art methods (Covington et al., 2016; Zhan et al., 2022; Sun et al., 2024; Lin et al., 2023) with our GR.
-For a fair comparison, similar model structures except for the output layers and corresponding loss functions are employed for these approaches.
-Details of the compared methods are provided in ”watch time prediction” of Sec.
-2.
-4.1.3. Metrics
-To evaluate the proposed method, we follow previous work (Lin et al., 2023; Sun et al., 2024) and utilize two performance metrics:
-  • Mean Average Error (MAE): It quantifies regression accuracy by averaging the absolute deviations between predicted values $\{\hat{y_{i}}\}_{i=1}^{N}$ and actual values $\{y_{i}\}_{i=1}^{N}$ as follows:
-    (10) $MAE=\frac{1}{N}\sum_{i=1}^{N}\left|\hat{y}_{i}-y_{i}\right|$
-  • XAUC (Zhan et al., 2022): This measure assesses the concordance between the predicted and actual ordering of watch time values.
-We uniformly sample pairs from the test set and calculate the XAUC by determining the percentile of samples that are correctly ordered.
-A higher XAUC indicates better model performance.
+Considering baseline methods compared in prior studies (Sun et al., 2024; Lin et al., 2023), we compare several state-of-the-art methods (Covington et al., 2016; Zhan et al., 2022; Zhao et al., 2024; Sun et al., 2024; Lin et al., 2023) with our GR.
+More details of the compared methods are provided in the supplementary material.
 4.2. Performance Comparison (RQ1)
-4.2.1. Offline
-We compare the proposed method GR with 5 existing models in terms of watch time and watch ratio, results are presented in Tab.
-2.
-Specifically, for watch time prediction, GR exhibits a 4.117% relative decrease in MAE and a 1.917% improvement in XAUC on the CIKM16 dataset.
-On the KuaiRec benchmark, significant improvements are achieved, with a 3.356% reduction in MAE and a 3.367% increase in XAUC relative to the second-best method.
-As for Indust dataset, GR exhibits a 3.629% relative decrease in MAE and a 1.001% improvement in XAUC compared to the CREAD.
-Regarding watch ratio predictions, all models exhibit significant performance gain due to the elimination of video duration bias.
-Despite this, GR maintains the best performance, boasting a 7.756% reduction in MAE and a 2.033% improvement in XAUC.
-The comprehensive improvements in both MAE and XAUC substantiate the superiority of the GR method.
-To confirm that the observed enhancements in the GR model are not solely attributable to increased model parameters, further comparisons with parameter-equivalent models can be found in the appendix A.2.1.
+4.2.1. Offline Evaluation
+Tab.
+1 shows the comparative results between GR and six baselines across three datasets.
+GR achieves consistent improvements in both MAE and XAUC metrics.
+For watch time prediction, GR maintains superior performance with 4.117% MAE reduction and a 1.917% XAUC improvement on CIKM16.
+On the KuaiRec, it significantly outperforms the second-best method with a 3.356% MAE reduction and 3.367% XAUC lift.
+As for Indust dataset, GR exhibits a 3.629% relative decrease in MAE and a 1.001% improvement in XAUC compared to the CREAD, which is a notable enhancement on a real-world business dataset.
+Regarding watch ratio predictions, while all models gain significantly from eliminating duration bias, GR maintains the best performance, boasting a 7.756% MAE reduction and a 2.033% XAUC improvement.
+The comprehensive improvements in both MAE and XAUC substantiate GR’s superiority.
+We also conduct experiments with parameter-equivalent models (see supplementary materials) to ensure the performance gains are not solely from increased model parameters.
 4.2.2. Online A/B Testing
-Besides offline experiments, we also conduct an online A/B test on Kuaishou App to demonstrate the real-world efficacy of our method.
-Considering that Kuaishou serves about 400 million users daily, doing experiments from 2.5% of traffic involves a huge population of more than 10 million users, which can yield highly reliable results.
-Most recommendation systems follow a two-stage framework: initially retrieving a set of candidate items and then selecting the top-ranked items in the ranking phase.
-Watch time prediction, a component of the ranking stage, plays a crucial role.
-Items with higher predicted watch times are more likely to be recommended.
-The online experiment have been launched on the system for six days and the results are in Tab.
-3.
-The metrics for online experiments include app usage time, average app usage per user, app launch counts and video consumption time (accumulated watch time).
-The results demonstrate that GR consistently boosts performance in watch time related metrics, with an improvement by 0.055% on app usage time, 0.072% on average app usage per user, and significant 0.129% on video consumption time, substantiating its potential to significantly enhance real-world user experiences.
-4.3. Vocabulary Construction Analysis (RQ2)
+[…]
+4.3. Underlying Reasons Analysis For Performance Gain (RQ2)
+We analyze model performance across ground truth (GT) watch time intervals on KuaiRec, where approximately 80% of videos have GT $\leq$10s.
+By splitting the range of watch time into 2-second long segments, Fig.
+5(a) shows that GR significantly outperforms CREAD and TPM for videos with short and medium watch times, with slightly lower performance only in the last ¿10s interval, where it lags behind TPM.
+For a more intuitive analysis, Fig.
+5(b-d) visualizes the distributions of Ground Truth (GT) watch times and the predictions generated by different methods, alongside their means and variances.
+Notably, the mean predicted watch time of GR closely aligns with the GT mean, whereas those of CREAD and TPM deviate significantly.
+Regarding variance, GR exhibits the largest spread, while CREAD shows the smallest.
+This corresponds visually to CREAD’s highly peaked distribution versus GR’s broader and flatter curve, suggesting GR’s capability to generate a more diverse and personalized set of predictions.
+Furthermore, GR is the only method that accurately predicts when GT is close to 0s, highlighting its flexibility afforded by its ability to output the ¡EOS¿ token in the first step.
+Fig.
+5(c) and (d) also visually confirm our hypothesis that CREAD and TPM tend to overestimate watch times, stemming from their rigid discretization structure where excessively large span values in tail buckets disproportionately amplify prediction errors, especially for videos with shorter watch times.
+As shown in Fig.
+5(d), the prediction distribution of TPM exhibits a notable skew towards higher values, attributable to the model’s tendency (observed during case analysis) to learn probabilities greater than 0.5 at the root node of its tree structure.
+This can result in an overall overestimation of the predicted outcomes, thereby explaining why GR’s performance is marginally surpassed by TPM in the ¿10s interval.
+However, given the characteristic long-tail distribution of real-world watch time data, the superior overall performance and distributional fidelity achieved by GR represent a favorable trade-off for this minor discrepancy in the high-value range.
+Figure 4. Token distribution comparison among vocabulary construction methods: (a) Manual, (b) Binary, (c) Dynamic.
+4.4. Vocabulary Construction Analysis (RQ3)
 Here we examine the effect of the vocabulary construction method.
@@
 In contrast, our dynamic quantile method achieves a more balanced distribution, further validating the efficacy of the proposed algorithm.
-Table 4. Comparison of vocabulary construction methods.
-Vocabulary design KuaiRec CIKM16
-[…]
-Manual 3.281 0.604 0.825 0.685
-Binary 3.268 0.605 0.821 0.687
-Dynamic quantile 3.196 0.614 0.815 0.691
-4.4. Generative Architecture Analysis (RQ3)
-To investigate the impact of different network architectures on prediction performance, we compare three popular sequence modeling structures: RNN, GRU, and LSTM, by substituting the decoder structure in the GR model.
-Results are presented in Fig.
-5, from which we can see that GR is not dependent on a specific network architecture and achieves desirable outcomes across various architectures.
-Among them, LSTM exhibits the best performance, likely because GR ultimately sums all predicted tokens to estimate watch time, resulting in strong dependencies between tokens in the sequence.
-LSTM is particularly adept at capturing these dependencies across different token positions.
-Figure 4. Token distribution comparison among vocabulary construction methods: (a) Manual, (b) Binary, (c) Dynamic.
-4.5. Effect of Curriculum Learning with Embedding Mixup (RQ4)
-Here, we investigate the impact of the proposed curriculum learning (CL) with embedding mixup (EM) strategy by designing several variants to perform comparisons from three perspectives.
-  • To study the contribution of CLEM, we design three ablation variants.
-w/o CLEM uses free running mode to directly feed $\bm{E}[\hat{s}_{i}^{t-1},:]$ to generate $\hat{s}_{i}^{t-1}$.
-EM with TF indicates embedding mixup strategy with full teacher forcing, i.e., with fixed sampling rate $p$ = 1.
-CL w/o EM removes embedding mixup strategy.
-  • To explore the effect of different dynamic decay strategy, we design linear decay and exponential decay instead of Eq. (7), thereby constructing two variants linear and exponential.
-  • To examine the effect of the curriculum strategy and free running, we construct two model variants with fixed sampling rates, where $p$ is set to 0.5 and 0, respectively.
-Table 5. Ablation study on the strategy of curriculum learning with embedding mixup.
+Figure 5.
+(a) Comparison of MAE on the KuaiRec dataset across videos with different watch time intervals.
+(b-d) The distribution comparison of predicted watch times among TPM, CREAD, and GR, compared to the Ground Truth (GT).
+Table 4. Ablation study on the strategy of curriculum learning (CL) with embedding mixup (EM).
 Method KuaiRec CIKM16
@@
 (h) ${p=0}$ 3.283 0.593 0.846 0.681
-The results in Tab.
-5 indicate that both strategies in CLEM enhance performance.
-Specifically, comparing row (a) with row (c), we can see that CL increases XAUC by 1.656% and reduces MAE by 1.38% on the KuaiRec dataset, while comparing row (a) with row (d) shows EM achieves even greater improvements, with 4.235% increase in XAUC and 4.853% reduction in MAE.
-Removing all the strategies results in substantial performance degradation for the GR model, with a loss of 5.137% in XAUC and 6.884% in MAE.
+4.5. Ablation study on Curriculum Learning with Embedding Mixup (RQ4)
+To systematically evaluate the proposed Curriculum Learning with Embedding Mixup (CLEM) framework, we conduct controlled ablation experiments across three dimensions:(1) component effectiveness, (2) scheduling sensitivity, and (3) nonlinear decay impact.
+The experimental variants a re designed as follows:
+  • Component Analysis:
+    • w/o CLEM: Vanilla training using direct feature projection $\bm{E}[\hat{s}_{i}^{t-1},:]\rightarrow\hat{s}_{i}^{t}$ without curriculum scheduling or mixup.
+    • EM with TF: Embedding mixup with full teacher forcing (fixed sampling rate $p=1$).
+    • CL w/o EM: Curriculum learning without embedding mixup regularization.
+  • Decay Strategy Comparison:
+    • Linear: Linear sampling rate decay $p_{t}=1-\tau t$.
+    • Exponential: Exponential decay $p_{t}=e^{-\tau t}$.
+  • Sampling Rate Impact:
+    • Fixed-0.5: Constant sampling rate $p=0.5$.
+    • Fixed-0: Pure free-running mode ($p=0$).
+As shown in Tab.
+4, the full CLEM framework (Row a) demonstrates significant improvements over baseline configurations.
+Compared to the non-curriculum variant (Row c), curriculum learning alone provides a 1.656% XAUC boost and 1.38% MAE reduction on the KuaiRec dataset.
+Embedding mixup contributes more substantially: disabling mixup (Row d) degrades XAUC by 4.235% and increases MAE by 4.853%, highlighting its critical regularization role.
 The sampling rate decay coefficients significantly impact both metrics.
@@
 These findings indicate that the CLEM strategy improves the model’s accuracy of watch time prediction.
-Figure 5. Performance comparison of different generative architectures on (a) KuaiRec and (b) CIKM16.
 4.6. Performance on LTV Prediction Task (RQ5)
-GR is a versatile regression framework.
-To explore its potential in other tasks, we conduct additional experiments on the Lifetime Value (LTV) prediction task.
-Following the settings  (Weng et al., 2024), we evaluate GR on Criteo-SSC [^4] https://ailab.criteo.com/criteo-sponsored-search-conversion-log-dataset/ and Kaggle [^5] https://www.kaggle.com/c/acquire-valued-shoppers-challenge, using MAE and Spearman rank correlation (Spearman’s $\rho$) as performance metrics.
-We consider all baseline methods compared in prior studies (Weng et al., 2024).
+GR is a generalized regression framework.
+To rigorously evaluate its cross-task generalization capability, we conduct extended experiments on the Lifetime Value (LTV) prediction task under identical experimental protocols as  (Weng et al., 2024).
+The evaluation employs two datasets: Criteo-SSC [^5] https://ailab.criteo.com/criteo-sponsored-search-conversion-log-dataset/ and Kaggle [^6] https://www.kaggle.com/c/acquire-valued-shoppers-challenge, with MAE and Spearman’s rank correlation (Spearman’s $\rho$) serving as performance metrics.
+All baseline implementations strictly adhere to the configurations documented in (Weng et al., 2024).
+As shown in Tab.
+5, GR achieves state-of-the-art performance with relative improvements of 17.66% in MAE and 20.79% in Spearman’s $\rho$ on Criteo-SSC over the previous best method OptDist (Weng et al., 2024).
+Notably, these baselines include task-specific architectures with dedicated LTV prediction modules.
+The consistent superiority of GR across both point estimation (MAE) and ranking correlation ($\rho$) metrics provides empirical evidence for its inherent robustness and domain-agnostic characteristics.
 Table 5. Performance comparison on LTV datasets.
@@
 GR (ours) 12.996 0.3026 67.035 0.5334
-Tab.
-6 presents the results of various methods in LTV prediction.
-These results show that our approach significantly outperforms the previous works across ordinal and numeric performance metrics, achieving 17.66% improvement of MAE and 20.79% improvement on Spearman’s $\rho$ on Criteo-SSC, compared with the state-of-the-art method OptDist (Weng et al., 2024).
-Given that the compared methods are specifically optimized for the LTV task, these findings demonstrate the robustness and extensibility of GR, underscoring its potential for application to other continuous value prediction tasks within the domain of recommendation systems.
 5. Conclusion
-This paper proposes a novel regression paradigm Generative Regression (GR) to accurately predict watch time, which addresses two key issues associated with existing methods.
-First, many ordinal regression methods struggle to accurately recover durations due to discretization, with performance heavily reliant on the chosen time binning strategy.
-Second, while ordinal regression implicitly constrains the probability distribution along the estimation path to exhibit a decreasing trend, existing methods have not fully leveraged this property.
-The generative regression approach builds upon autoregressive modeling and offers a promising exploration space.
-We also introduce embedding mixups and curriculum learning during training to accelerate model convergence and enhance performance.
-Extensive offline and online experiments show that GR significantly outperforms the state-of-the-art models.
-Additionally, we investigate its applicability to lifetime value (LTV) prediction, where our generative regression method also surpasses the SOTA models, highlighting its potential as a novel and effective solution to general regression challenges.
-Appendix A Appendices
-A.1. Experimental Settings
-Unless otherwise stated, we adopt the following training protocol.
-The hyperparameters $\lambda$ in equation Eq.
-(6) are deliberately set to 10.
-Our GR architecture comprises a two-layer Multi-Layer Perceptron (MLP) for the encoder and a single-layer Long Short-Term Memory (LSTM) network for the decoder.
-We employ a dropout rate of 0.3 to prevent overfitting and set the learning rate to 5e-4.
-As for vocabulary construction in Algorithm.
-1, the decay rate $\alpha$ is set to 0.99, precision threshold $\epsilon$ is set to 1e-6, start percentile $q_{start}$ and end percentile $q_{e}nd$ are respectively set to 0.9999 and 0.90.
-The $\omega$ in Eq.
-(7) is set to 17000.
-In Eq.
-(8), the window size $n_{w}$ is set to half the length of the vocabulary.
-We train for 20 epochs in total with a batch size of 1024, using the Adam optimizer with default parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ to minimize the loss function.
-Table 7. The impact of model parameters on the performance between different approaches on KuaiRec and CIKM16.
-Method Parameters KuaiRec CIKM16
-MAE $\downarrow$ XAUC $\uparrow$ MAE $\downarrow$ XAUC $\uparrow$
-VR 0.86M 7.634 0.504 1.039 0.641
-TPM 0.86M 3.456 0.571 0.850 0.676
-CREAD 0.86M 3.307 0.594 0.865 0.678
-VR-large 4.34M 7.556 0.521 0.998 0.651
-TPM-large 4.34M 3.432 0.577 0.843 0.682
-CREAD-large 4.34M 3.293 0.599 0.857 0.680
-GR (ours) 4.18M 3.196 0.614 0.815 0.691
-A.2. Additional Ablation Study Results
-A.2.1. Model Parameters
-To ascertain that the enhancement observed in the GR model is not merely a consequence of an increase in model parameters, we conducted a comparative analysis, as presented in Tab.
-7.
-Specifically, we standardized the parameters of the SOTA models, CREAD (Sun et al., 2024) and TPM (Lin et al., 2023), to a uniform level.
-The results indicate that, when evaluated with an equivalent parameter scale, the GR model consistently surpasses the previous SOTA methods across a range of metrics.
-A.2.2. Hyperparameter for Vocabulary Construction
-We conduct an ablation study to investigate the effects of the decay rate $\alpha$ and end percentile $q_{end}$ in Algorithm.
-1 on the performance of GR.
-All experiments maintain a constant start percentile $q_{start}$ set at 0.9999 and the results are present in Tab.
-8.
-Under the same end percentile $q_{end}$, models with a decay rate of 0.99 consistently outperform those with a decay rate of 0.98.
-Within identical settings for $\alpha$, the configurations with $q_{end}=0.9$ yield the best performance results, followed by $q_{end}=0.85$.
-Figure 6.
-(a) Dynamic probability across different epochs for varying parameter $\omega$.
-The curves show the decrease in sampling rate over time.
-(b) The impact of $\omega$ on MAE and XAUC for the KuaiRec dataset.
-A.2.3. Hyperparameter for Curriculum Learning (CL)
-Here, we check the effects of the hyperparameters $\omega$ in Eq.
-(7) which influences the shape of descent curve to ensure a seamless transition from higher to lower values.
-As shown in Fig.
-6(a), four distinct values of $\omega$, i.e., 4600, 8500, 13000, and 17000, are chosen to control the dynamic probability, gradually reducing it to zero at the 5th, 10th, 15th, and 20th epoch, respectively.
-The results are presented in Fig.
-6(b).
-We can see that $\omega=13000$, i.e., dynamic probability approaching zero around the 15th epoch, performs better than the other settings.
-$\omega=4600$ (dynamic probability approaching zero around the 5th epoch) results in a premature reduction in sampling probability, leading to the worst performance.
-Meanwhile, the performance with $\omega=17000$, while slightly inferior to that of $\omega=13000$, still surpasses the outcomes of strategies where the probability reduces to zero too prematurely.
-Therefore, judiciously setting the $\omega$ parameter, and learning to predict with both ground truth and previous prediction as input can effectively resolve inconsistencies between training and inference.
-Figure 7. Performance on KuaiRec dataset with different hyperparameter settings in vocabulary construction.
-MAE $\downarrow$ XAUC $\uparrow$
-$\alpha=0.98$ $\small q_{end}=0.95$ 3.282 0.574
-$\small q_{end}=0.90$ 3.216 0.601
-$\small q_{end}=0.85$ 3.227 0.589
-$\small q_{end}=0.80$ 3.256 0.578
-$\alpha=0.99$ $\small q_{end}=0.95$ 3.253 0.590
-$\small q_{end}=0.90$ 3.196 0.614
-$\small q_{end}=0.85$ 3.205 0.603
-$\small q_{end}=0.80$ 3.231 0.595
-Figure 8. The impact of window size $n_{w}$ on MAE and XAUC for KuaiRec dataset.
-A.2.4. Hyperparameter for Embedding Mixup (EM)
-Here, we check the effects of the window size $w_{n}$ in Eq.
-(8) which control the the fixed size of window to merge representations of tokens within it, preventing significant errors.
-As shown in Fig.
-8, we select six percentages of the vocabulary size to determine the window size values for our ablation experiments on the Kuairec dataset.
-The horizontal axis represents these percentages (actual window sizes), and we also expand the window size to include weighting the entire vocabulary (denoted as ”all”).
-The results show that as the percentage increases, performance gradually improves, reaching its peak at 50%, and further increases in the window size lead to a slight performance decline.
-A smaller window size may fail to merge meaningful token representations, resulting in suboptimal performance.
-Conversely, a larger window size may introduce noise by incorporating less relevant tokens, which can degrade performance.
+This paper proposes a novel regression paradigm Generative Regression (GR) to accurately predict watch time, which addresses two key issues associated with existing ordinal regression (OR) methods.
+First, OR struggles to accurately recover watch times due to discretization, with performance heavily reliant on the chosen time-binning strategy.
+Second, while OR implicitly constrains the probability distribution along the estimation path to exhibit a decreasing trend, existing methods have not fully leveraged this property.
+GR builds upon autoregressive modeling and offers a promising exploration space.
+We also introduce embedding mixups and curriculum learning during training to accelerate model convergence.
+Extensive online and offline experiments show that GR significantly outperforms the SOTA models.
+Additionally, our GR also surpasses the SOTA models in lifetime value (LTV) prediction, highlighting its potential as an effective general regression solution.